### v1

# DisastIR: A Comprehensive Information Retrieval Benchmark for Disaster Management

Abstract
Effective disaster management requires timely access to accurate and contextually relevant information. Existing Information Retrieval (IR) benchmarks, however, focus primarily on general or specialized domains, such as medicine or finance, neglecting the unique linguistic complexity and diverse information needs encountered in disaster management scenarios. To bridge this gap, we introduce DisastIR, the first comprehensive IR evaluation benchmark specifically tailored for disaster management. DisastIR comprises 9,600 diverse user queries and more than 1.3 million labeled query-passage pairs, covering 48 distinct retrieval tasks derived from six search intents and eight general disaster categories that include 301 specific event types. Our evaluations of 30 state-of-the-art retrieval models demonstrate significant performance variances across tasks, with no single model excelling universally. Furthermore, comparative analyses reveal significant performance gaps between general-domain and disaster management-specific tasks, highlighting the necessity of disaster management-specific benchmarks for guiding IR model selection to support effective decision-making in disaster management scenarios. All source codes and DisastIR are available at this repository.
DisastIR: A Comprehensive Information Retrieval Benchmark for
Disaster Management
Kai Yin1  Xiangjue Dong1  Chengkai Liu1  Lipai Huang1 Yiming Xiao1Zhewei Liu2Ali Mostafavi1James Caverlee1 1Texas A&M University  2University of Toronto {kai_yin,xj.dong,liuchengkai,lipai.huang,yxiao,mostafavi,caverlee}@tamu.edu zwei.liu@utoronto.ca
1 Introduction
Natural disasters and technological crises cause severe threats to human lives, infrastructure, and the environment, necessitating timely and effective management responses (Dong et al., 2020; Yin et al., 2023; Liu et al., 2024). In such critical scenarios, stakeholders, including emergency responders, government agencies, and the general public, require rapid access to reliable and contextually relevant information to make informed decisions (Jayawardene et al., 2021; Abbas and Miller, 2025). Information Retrieval (IR) systems thus play a critical role in disaster management, where rapid, accurate access to relevant information can significantly impact emergency response outcomes and decision-making efficacy (Basu and Das, 2020; Kumar et al., 2023; Langford and Gulla, 2024).
Figure 1: Examples of user queries across diverse search intents and event types during disaster management.
Figure 2: Proposed framework to develop DisastIR from scratch.
Information needs during real-world disasters are highly diverse (Figure 1), including intents such as question answering, rumor verification, social media monitoring, and evidence retrieval (Purohit et al., 2014; Imran et al., 2015; Zubiaga et al., 2018). These varied intents require tailored retrieval behavior (Asai et al., 2022; Su et al., 2022; Lee et al., 2024b) and understanding of “relevance” (Dai et al., 2022). In addition, different types of disasters (Figure 1), such as geohazards, biological threats, and technological failures, differ significantly in terminology, phrasing, and discourse styles (Andharia, 2020; UNDRR, 2020; Bromhead, 2021). This complexity presents significant challenges for retrieval systems aiming to serve real-world disaster response scenarios.
However, existing retrieval benchmarks primarily target general-domain tasks, such as BEIR (Thakur et al., 2021), or focus on specific domains like medicine (Wang et al., 2024a) and finance (Tang et al., 2024). They are not designed to reflect the search task diversity and domain-specific demands of disaster management scenarios. As a result, current IR evaluation benchmarks offer limited guidance for selecting retrieval models in disaster management applications.
To address this gap, we present DisastIR, the first comprehensive IR benchmark tailored to disaster management. DisastIR evaluates retrieval models across 48 distinct tasks, defined by combinations of six real-world search intents and eight general disaster event types, covering a total of 301 specific event types (see Section 3.2).
DisastIR is built on a systematically constructed disaster management-specific corpus, developed through extensive web crawling, semantic chunking, and deduplication (Section 3.3). To simulate realistic information needs, we use a large language model (LLM) [^1] The LLM used in this work is GPT-4o-mini. to generate diverse, contextually grounded user queries (Section 3.4). Candidate passages are aggregated from multiple state-of-the-art (SOTA) retrieval models (Section 3.5), and query-passage pairs are annotated using LLMs with three different designed prompts whose outputs are ensembled for robust relevance labeling (Section 3.6).
To ensure annotation quality and evaluation reliability, we validate LLM-generated relevance labels against human annotations, observing substantial agreement (average Cohen’s kappa = 0.77; see Section 4.2). We also compare LLM-generated and human-written queries across all 48 tasks (Section 4.3) and find highly consistent evaluation results (Kendall’s $\tau$ = 0.93), supporting the use of synthetic queries and relevance labels in DisastIR.
Using DisastIR, we benchmark 30 open-source retrieval models of varying sizes, architectures, and backbones under both exact and approximate nearest neighbor (ANN) search settings (Section 5). Our results show that no single model consistently outperforms others across all disaster management-related retrieval tasks (Section 6.2). We also observe substantial performance gaps between general-domain benchmarks (e.g., MTEB (Muennighoff et al., 2022)) and DisastIR (Section 6.3), highlighting the need for a domain-specific benchmark to guide reliable and effective retrieval model selection in disaster management scenarios.
The contributions of this work are as follows:
  • (1)
    We release DisastIR, the first IR benchmark tailored to disaster management. It includes a systematically constructed evaluation corpus of 239,704 passages and 9,600 user queries, with over 1.3 million annotated query-passage pairs across 48 retrieval tasks spanning diverse search intents and disaster event types.
  • (2)
    We conduct a comprehensive evaluation of 30 open-source retrieval models under both exact and ANN search settings, offering practical guidance for model selection based on task requirements and computational constraints in disaster management scenarios.
  • (3)
    We empirically demonstrate substantial performance gaps between general-domain and disaster management-specific retrieval, underscoring the necessity of disaster management-specific IR evaluation benchmarks.
2 Related work
Existing IR benchmarks target mainly general-purpose or specialized domains, such as medicine and finance. BEIR (Thakur et al., 2021) evaluates zero-shot retrieval models across 18 tasks, such as fact verification, QA, and scientific document ranking. Instruction-based benchmarks like FollowIR (Weller et al., 2024), InstructIR (Oh et al., 2023), and MAIR (Zhang et al., 2024b) reformulate IR tasks using natural language instructions. Some domain-specific IR benchmarks, such as MIRAGE (Wang et al., 2024a) and FinMTEB (Tang et al., 2024), focus on biomedical and financial domains. While effective in their respective domains, they fail to capture the linguistic and contextual patterns in disaster management areas.
Despite the critical role of information retrieval in disaster management, existing benchmarks are limited in scope, scale, and task diversity. Prior datasets—such as the FIRE IRMiDis track (Basu et al., 2017) and event-specific corpora from disasters in Nepal, Italy, and Indonesia (Khosla et al., 2017; Basu and Das, 2019; Kumar et al., 2023)—primarily focus on Twitter microblogs, targeting short-text retrieval or keyword matching with narrow task coverage. Case-based systems like Langford and Gulla (2024) use proprietary data for concept-based retrieval in search and rescue planning. These benchmarks typically rely on single-source or scenario-specific data and lack support for realistic, multi-intent retrieval. In contrast, DisastIR provides a large-scale, multi-intent, and multi-source benchmark covering diverse disaster types and information needs, enabling comprehensive evaluation in real-world contexts.
3 DisastIR: Disaster Management Information Retrieval Benchmark
3.1 Overview
The construction of DisastIR follows a four-stage pipeline, as illustrated in Figure 2: (1) disaster management corpus construction, (2) user query generation, (3) candidate pool development, and (4) relevance labeling. DisastIR is built upon a large-scale, high-quality corpus of disaster management-related passages covering diverse event types. User queries are generated by prompting an LLM with these domain passages as context, targeting different search intents. Relevance scores for each query-passage pair are then assigned by the LLM.
3.2 Evaluation Task
To evaluate how well retrieval models address diverse user intents and disaster contexts, DisastIR defines six search intents and eight general disaster event types, resulting in 48 distinct retrieval tasks.
Specifically, 301 specific event types are identified spanning eight general categories: Biological (Bio), Chemical (Chem), Environmental (Env), Extraterrestrial (Extra), Geohazard (Geo), Meteorological & Hydrological (MH), Societal (Soc), and Technological (Tech) (UNDRR, 2020). See Figure 1 for examples of specific event types belonging to each general disaster event type.
Six distinct search intents are included, inspired by prior benchmarks such as BEIR (Thakur et al., 2021), BERRI (Asai et al., 2022), MEDI (Su et al., 2022), and MAIR (Sun et al., 2024): question-answer (QA) retrieval, Twitter retrieval, Fact Checking (FC) retrieval, Natural Language Inference (NLI) retrieval, and Semantic Textual Similarity (STS) Retrieval. For QA, we further distinguish between retrieving relevant passages (QA) and retrieving relevant documents (QAdoc), following common practice in prior work (Kwiatkowski et al., 2019; Khashabi et al., 2021; Xu et al., 2024). [^2] A passage refers to a single chunk with limited token length, while a document denotes a full source file, which may be segmented into multiple passages. Due to token limitations in many retrieval models – especially encoder-based ones – it is often infeasible to encode full documents directly. To address this, we prompt an LLM to summarize each document and include the summary in the corpus as a proxy for the original document.
3.3 Domain knowledge corpus construction
To construct the domain knowledge corpus, we perform a large-scale web crawling using 301 disaster event types as search queries, collecting domain-specific PDF documents from publicly available sources. A structured pipeline is then applied to convert raw PDFs into clean, retrieval-ready passages: (1) exact-URL deduplication, (2) text extraction and preprocessing, (3) document-level near-duplicate removal using locality-sensitive hashing (LSH), (4) semantic chunking, and (5) embedding-based near-duplicate filtering. The full pipeline is described in Appendix A.
3.4 User Query Generation
A key challenge in constructing domain-specific IR evaluation datasets is generating user queries that reflect real information needs (Rahmani et al., 2024a). With the advent of LLMs, it is now feasible to synthesize high-quality, diverse, and contextually grounded queries by prompting models with domain-specific passages (Alaofi et al., 2023; Rajapakse and de Rijke, 2023; Rahmani et al., 2024a).
In this work, we propose a two-stage few-shot prompting strategy to generate user queries based on disaster management passages. In the first stage, an LLM is prompted to brainstorm diverse information need statements grounded in the content of the given passage. In the second stage, given a randomly selected information need and the associated passage, the LLM generates a user query and a directly relevant passage as shown below:
$LLM_{query}\bigl{(}\underbrace{LLM_{info}(P_{IN},P_{seed})}_{\displaystyle\text{information need}},\;P_{QG},\;P_{seed}\bigr{)}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\quad\longrightarrow\;(q,\;psg)~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}(1)$
where $LLM_{info}$ and $LLM_{query}$ are LLMs prompted to generate retrieval information needs statements and the query-passage pair respectively, $P_{IN}$ and $P_{QG}$ are prompts for information needs and query generation, $P_{seed}$ is the domain passage, $q$ is the synthesized user query, and $psg$ is the corresponding relevant passage.
To ensure generated queries align with the core characteristics and objectives of each search intent, we design intent-specific prompts for both stages of query generation. The full prompt templates for each intent are provided in Appendix B.
For each search task, we generate 200 unique user queries by prompting an LLM with randomly sampled domain-specific passages, resulting in 9,600 queries. The final corpus combines disaster management-related passages from Section 3.3 with generated passages to reflect various search intents. Some tasks, such as Twitter, NLI, and FC retrieval, require passage types with distinct styles and semantics. Including generated passages ensures the corpus can support realistic evaluation across diverse retrieval scenarios. [^3] Relevance scores of query-generated passage pairs are also evaluated instead of directly giving them the highest relevance score.
QA QAdoc Twitter FC NLI STS
Bio 26651 (133.3) 25335 (126.7) 35182 (175.9) 23987 (119.9) 25896 (129.5) 27065 (135.3)
Chem 26885 (134.4) 26032 (130.2) 34186 (170.9) 24592 (123.0) 27856 (139.3) 26787 (133.9)
Env 26685 (133.4) 25930 (129.7) 33243 (166.2) 25805 (129.0) 25207 (126.0) 27048 (135.2)
Extra 26807 (134.0) 25598 (128.0) 33202 (166.0) 24363 (121.8) 26399 (132.0) 27313 (136.6)
Geo 27140 (135.7) 26573 (132.9) 35503 (177.5) 27864 (139.3) 28210 (141.1) 29816 (149.1)
MH 28422 (142.1) 27256 (136.3) 33924 (169.6) 26670 (133.4) 27052 (135.3) 28702 (143.5)
Soc 27116 (135.6) 23353 (116.8) 33834 (169.2) 27850 (139.3) 26997 (135.0) 27074 (135.4)
Tech 28044 (140.2) 27071 (135.4) 33388 (166.9) 26759 (133.8) 28394 (142.0) 26920 (134.6)
Table 1: Number of labeled query-passage pairs and pairs per query (in parentheses) of each search task in DisastIR.
3.5 Assessment Candidate Pool Development
Given the large size of the corpus, annotating all possible (query, passage) pairs is impossible (Thakur et al., 2021). Following prior work, we construct a candidate pool for each query using existing retrieval models. Inspired by TREC’s standard practice, where top-ranked passages from multiple systems are aggregated to form the candidate set, we adopt a similar strategy in DisastIR.
Specifically, for each query, we collect the top $10$ retrieved passages from 30 retrieval models under two retrieval settings: exact and ANN search settings (detailed in Section 5). These models also serve as baselines for performance evaluation, following practices in recent work (Rahmani et al., 2024b; Wang et al., 2024b). The candidate pool for each query is formed by taking the union of passages retrieved under both settings.
3.6 Relevance Labeling
Once query-passage pairs are prepared, we annotate them using an LLM. Recent studies have shown that LLMs can reliably produce relevance judgments that align closely with human annotations (Rahmani et al., 2024a, b, 2025; Wang et al., 2024b). Furthermore, Wang et al. (2024b); Rahmani et al. (2024c) demonstrate that ensembling relevance scores from multiple prompts or LLMs yields more robust and calibrated annotations.
To this end, we design three diverse prompts for each search intent and use a single LLM to generate relevance scores. The prompts, inspired by Thomas et al. (2024); Farzi and Dietz (2024); Rahmani et al. (2025), are: (1) zero-shot direct scoring—a single-pass judgment; (2) chain-of-thought reasoning—a multi-step prompt mimicking human-style reasoning; and (3) multi-dimensional attribute scoring—relevance decomposed into interpretable sub-criteria. For each search intent, relevance is defined to align with its specific objectives, reflecting the varying interpretations of “relevance” across different task types (Dai et al., 2022). Full prompt templates are provided in Appendix C.
Relevance scores are assigned on a 4-point scale (0 to 3) for all intents, except STS, which follows a 6-level scale as in Agirre et al. (2013); Cer et al. (2017). The final score for each pair is computed by averaging scores from three prompts.
Count Avg Median Min Max
Query 9,600 33.75 19 2 281
Passage 239,704 197.17 224 6 2,536
Table 2: Statistics of number of query and passage and their token lengths. Tokenization is based on the cl100k_base tokenizer (used in GPT-4 / GPT-3.5).
Model Param Size Exact $\uparrow$ Ex. ANN Drop
Name Size Bin QA QAdoc TW FC NLI STS Avg Avg (%)
inf-retriever-v1 7B XL 74.23 67.82 69.33 68.91 53.10 77.70 68.52 66.90 2.36
SFR-Embedding-Mistral 7B XL 71.76 67.58 70.42 70.62 50.86 73.61 67.47 66.75 1.08
inf-retriever-v1-1.5b 1.5B XL 70.64 64.42 66.11 66.53 53.86 76.19 66.29 65.54 1.13
NV-Embed-v2 7B XL 74.77 69.74 43.18 68.64 58.73 77.01 65.34 64.45 1.36
multilingual-e5-large 560M Large 67.29 64.25 63.75 60.26 51.02 75.06 63.61 62.08 2.41
multilingual-e5-large-instruct 560M Large 68.39 64.90 63.24 67.21 49.38 64.31 62.91 62.06 1.34
e5-small-v2 33M Small 65.87 63.00 60.89 62.04 47.09 74.83 62.29 60.91 2.22
e5-base-v2 109M Medium 65.77 63.07 58.37 62.28 45.54 74.64 61.61 60.23 2.25
e5-large-v2 335M Large 60.15 63.42 56.20 62.29 50.96 74.99 61.34 60.20 1.85
NV-Embed-v1 7B XL 68.36 63.02 56.84 60.04 48.31 67.86 60.74 59.24 2.47
granite-embedding-125m 125M Medium 64.90 61.04 47.14 62.78 48.04 71.94 59.31 58.67 1.06
llmrails-ember-v1 335M Large 64.82 63.34 45.93 61.06 44.56 73.99 58.95 58.35 1.02
arctic-embed-m-v2.0 305M Medium 61.44 62.50 47.80 58.04 42.34 65.28 56.23 54.75 2.64
mxbai-embed-large-v1 335M Large 64.57 62.96 40.58 58.47 40.22 68.81 55.93 55.42 0.92
gte-base-en-v1.5 137M Medium 60.66 56.06 47.06 52.49 39.88 71.27 54.57 53.71 1.57
gte-large-en-v1.5 434M Large 67.70 58.55 40.31 53.09 34.81 67.30 53.63 53.09 1.01
snowflake-arctic-embed-l-v2.0 568M Large 55.34 59.43 38.75 60.42 41.19 63.38 53.09 51.90 2.23
bge-base-en-v1.5 109M Medium 49.67 53.97 47.03 58.47 35.58 64.43 51.52 50.71 1.57
bge-large-en-v1.5 335M Large 57.02 54.72 32.71 55.22 35.19 65.14 50.00 48.99 2.03
gte-Qwen2-7B-instruct 7B XL 62.47 39.12 41.40 29.98 47.28 70.98 48.54 47.07 3.03
bge-small-en-v1.5 33M Small 47.38 45.51 27.82 52.13 27.20 56.86 42.82 42.07 1.75
snowflake-arctic-embed-m-v1.5 109M Medium 25.73 30.56 18.31 48.24 43.04 64.95 38.47 36.25 5.79
snowflake-arctic-embed-s 33M Small 36.05 27.46 18.17 42.52 37.27 66.59 38.01 33.59 11.63
snowflake-arctic-embed-l 335M Large 40.82 30.41 15.32 32.70 34.62 56.82 35.11 31.35 10.72
Linq-Embed-Mistral 7B XL 35.61 31.41 26.55 39.46 27.30 45.69 34.34 33.52 2.38
thenlper-gte-base 109M Medium 9.22 5.35 38.54 60.80 42.85 46.64 33.90 31.78 6.24
snowflake-arctic-embed-m 109M Medium 33.35 14.25 8.64 35.30 38.93 56.88 31.23 28.83 7.67
all-mpnet-base-v2 109M Medium 15.06 9.77 16.17 46.30 27.38 37.23 25.32 25.05 1.08
e5-mistral-7b-instruct 7B XL 21.65 19.51 19.48 31.04 20.39 39.57 25.27 24.05 4.82
gte-Qwen2-1.5B-instruct 1.5B XL 13.98 22.21 19.61 23.90 18.00 31.20 21.48 21.27 0.98
Table 3: Performances of 30 evaluated IR models in DisastIR. Models are ranked by their overall performance under exact search (highest to lowest) in DisastIR. “Size Bin” indicates its model parameter size bin category (small, medium, large, and extra large as defined in Appendix G). “TW” represents Twitter. Overall performance across all queries under exact and ANN search are in “Ex. Avg” and “ANN Avg” columns. “Drop” shows the percentage decrease from exact to ANN average scores. Bold indicates the highest value, and underline indicates the second-highest.
4 DisastIR Benchmark Analysis
4.1 Query and Passage Characteristics
Query and Passage Lengths.
As shown in Table 2, the average query length is 33.75 tokens, with a median of 19, and a long tail extending to 281 tokens. This variation reflects the diversity of search intents, from short entity-style queries to detailed information needs typical in real-world disaster management scenarios. Passages are much longer on average (197.17 tokens), with a median of 224, and some exceeding 2,500 tokens. This wide distribution captures the diversity of disaster management-related texts, including both brief updates and detailed descriptions like event summaries or emergency protocols.
Labeled Query-Passage Pairs.
Table 1 summarizes the distribution of labeled query-passage pairs. In total, we obtained 1,341,986 labeled pairs, with each query linked to an average of 140 passages.
As shown in Table 1, Twitter-related search tasks tend to have a higher average number of query-passage pairs per query. The candidate pool for each query is built by merging the top 10 passages retrieved by 30 different models. This larger pool in Twitter tasks suggests greater divergence in model outputs, indicating lower agreement among retrieval models when ranking passages in social media contexts within disaster management scenarios. Additional analyses of labeled query-passage pairs are provided in Appendix D.
4.2 LLM-based vs. Human Labeling
Since relevance scores in DisastIR are judged by LLM, it is vital to evaluate their consistency with human annotations. Thus, we construct the LVHL dataset (LLM-based Vs. Human Labeling) by sampling disaster management-related query-passage pairs with human-labeled relevance scores from several open-source datasets. MS MARCO (Bajaj et al., 2016) and TriviaQA (Joshi et al., 2017) are for QA, ALLNLI (sentence-transformers, 2021) and XNLI (Conneau et al., 2018) for NLI, Climate-Fever (Diggelmann et al., 2021) for FC, and STSB (Cer et al., 2017) for STS. Appendix E provides details on the construction of LVHL.
The LLM-based relevance scores for each query-passage pair in LVHL are computed as described in Section 3.6. Since most human-annotated relevance scores in LVHL are binary, we follow Wang et al. (2024b) and binarize the LLM scores into two levels: relevant (score > 0) and not relevant (score = 0), to enable meaningful comparison.
To assess agreement between LLM-based and human relevance labeling, we compute Cohen’s kappa for each search intent. All datasets yield kappa scores above 0.6 (Figure 6), with an average of 0.77, indicating substantial agreement. These suggest that LLM-generated relevance scores align well with human judgments and can reliably substitute for manual annotation in DisastIR.
4.3 LLM vs. Human-generated User Query
To evaluate whether LLM-generated queries can serve as a reliable alternative to human-authored ones for retrieval benchmarking, we construct LVHQ (LLM Vs. Human-generated Query), a comparison set spanning all 48 retrieval tasks. For each task, both an LLM-generated and a human-written query are created based on the same domain passage. All query-passage pairs are annotated using the same method as in DisastIR. Appendix F provides full details on the construction of LVHQ.
We evaluate all selected baseline models using LVHQ under exact search for both human- and LLM-generated queries (see Section 5 for evaluation setup up). Model performance, measured by NDCG@10, shows highly consistent results across the two query types, with a Kendall’s $\tau$ of 0.9264, indicating strong agreement in model evaluations.
5 Experimental Setup
5.1 Models
DisastIR is adopted to comprehensively evaluate open-source IR models and support the selection of suitable IR models for real-world disaster management applications. Models are chosen based on three criteria: (1) strong performance on the MTEB retrieval benchmark; (2) inclusion in widely adopted embedding families such as BGE (Chen et al., 2024; Xiao et al., 2024), E5 (Wang et al., 2022, 2023), Snowflake Arctic (Merrick, 2024), and GTE (Li et al., 2023; Zhang et al., 2024a), which are commonly used as baselines and in downstream IR tasks (Sun et al., 2024; Xu et al., 2024; Lee et al., 2024b, a; Cao, 2025; Park et al., 2025); and (3) coverage of different types of search tasks adapted from Wang et al. (2024b).
We select 30 models with parameter sizes ranging from 33 million to 7 billion. Detailed descriptions of these models and their implementations are provided in Appendix G.
Figure 3: Distribution of evaluated models’ performances across all 48 tasks. The full name of each model in the X axis is listed in Model Name column in Table 3.
5.2 Evaluation
We evaluate model performance under two retrieval settings – exact and ANN – using Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) as the primary metric, consistent with prior works.
(1) Exact Brute-force Retrieval.
Following prior work such as BEIR (Thakur et al., 2021), InstructIR (Oh et al., 2023), FollowIR (Weller et al., 2024), and MAIR (Zhang et al., 2024b), we compute similarity scores between each user query and all passages in the corpus, retrieving the top $k$ most similar ones. This setting reflects model performance under ideal retrieval conditions.
(2) Approximate Nearest Neighbor (ANN) Retrieval.
For large-scale corpora, brute-force retrieval is computationally infeasible. A common solution is a multi-stage architecture, where an ANN search retrieves a candidate set of passages, which are then re-ranked for final output (Tu et al., 2020; Macdonald and Tonellotto, 2021). To reflect real-world large-scale disaster information retrieval scenarios, we also evaluate model performance during the candidate generation stage using ANN search. We adopt Usearch (Vardanian, 2023), a high-performance, memory-efficient library based on the HNSW (hierarchical navigable small world) algorithm (Malkov and Yashunin, 2018), to retrieve top $k$ passages per query using precomputed embeddings. For fair comparison, $k$ is set to match the value used in exact search.
Figure 4: Comparison between DisastIR and MTEB model rankings. Legend shapes indicate the model size bin: $\blacklozenge$ XL, $\blacktriangle$ Large, $\blacksquare$ Medium, $\bullet$ Small.
6 Evaluation Results
6.1 Overall Performance
Table 3 summarizes the overall performance of all 30 evaluated models across all queries in DisastIR, with detailed results for each search task provided in Appendix H. The inf-retriever-v1 model ranks highest in both exact and ANN search settings, followed closely by SFR-Embedding-Mistral (1.52% and 0.22% lower, respectively). Among non-XL models, multilingual-e5-large performs best, reaching 92.83% and 92.79% of the top model’s performance. Notably, the lightweight e5-small-v2 model (33M parameters) achieves 90.91% and 91.05% of the top model’s performance, despite being 212 times smaller in size.
The Snowflake-arctic-embed-s model shows the largest performance drop (11.63%) under ANN search compared to exact search (Table 3). Most models exhibit drops within 5%; only five exceeded this margin, four of which belong to snowflake family, indicating strong robustness when switching from exact to ANN search in DisastIR. Among the top-performing models analyzed above, SFR-Embedding-Mistral has the smallest drop (1.08%), while multilingual-e5-large has the largest (2.41%). All subsequent analyses are based on exact search; analyses under ANN search can be conducted similarly.
Figure 5: Best-performing models in each search task.
6.2 Performance across all 48 Tasks
Figure 3 presents the performance distribution of all evaluated models across all 48 search tasks. While inf-retriever-v1 achieves the highest average NDCG@10, its median performance is lower than that of SFR-Embedding-Mistral, and it exhibits greater variability across tasks, as reflected by a larger interquartile range (IQR). This suggests that inf-retriever-v1 is less stable across diverse search tasks in DisastIR.
As shown in Figure 5, no single model consistently outperforms others across all 48 tasks. Instead, top performance is distributed among three models: inf-retriever-v1, SFR-Embedding-Mistral, and NV-Embed-v2. inf-retriever-v1 achieves the best results in only nine tasks, primarily in STS. This highlights the complexity and diversity of disaster management-related retrieval tasks and reinforces the need for domain-specific IR models in real-world disaster management scenarios. Appendix I provides additional analyses of model performance across 48 tasks.
6.3 Comparison with General Domain
Figure 4 compares model rankings in DisastIR and MTEB. Ranking value of each model is based on overall performance in DisastIR and official retrieval scores from the MTEB English leaderboard. The Spearman correlation between the two rankings is 0.225 ($p=0.251$), indicating no significant correlation. This suggests that strong performance on general-domain benchmarks does not guarantee effectiveness in disaster management-related retrieval. For example, models like Linq-Embed-Mistral and snowflake-arctic-embed-l perform well in MTEB but poorly in DisastIR, while models from the E5 family show the opposite trend.
Although inf-retriever-v1 ranks highest in both DisastIR and MTEB, detailed analysis in Section 6.2 show that it achieves top performance in fewer than 20% of DisastIR tasks, mostly those related to STS. Furthermore, when computational resources are limited and large models are impractical to serve, relying solely on MTEB rankings for model selection, such as choosing snowflake-arctic-embed-l, may fail to retrieve critical or relevant content. These discrepancies underscore the necessity of a domain-specific benchmark like DisastIR to guide retrieval model selection across different disaster management-related search tasks.
7 Conclusion
In this work, we introduce and publicly release DisastIR, the first comprehensive retrieval benchmark for evaluating model performance in disaster management contexts. DisastIR consists of 9,600 user queries and more than 1.3 million labeled query-passage pairs, spanning 48 retrieval tasks defined by six search intents and eight general disaster event types, covering 301 specific event types.
Using DisastIR, we evaluate 30 SOTA open-source retrieval models under both exact and ANN search settings. Our findings provide practical guidance for selecting appropriate IR models based on task type and computational constraints, supporting timely and effective access to critical information in disaster management scenarios.
Limitations
While DisastIR represents a significant step toward domain-specific evaluation in disaster information retrieval, several aspects merit further enhancement. DisastIR currently focuses on English-language resources. Expanding DisastIR to multilingual settings would enable broader applicability. Furthermore, tables and figures in domain-specific PDF files may contain useful domain knowledge. Further study could consider extracting this critical information for evaluation set development.
Ethics Statement
DisastIR is designed to support disaster management by improving the evaluation and selection of retrieval models. All data used in the benchmark are sourced from publicly available materials, and no personally identifiable information is included. All contents generated by LLMs are evaluated by a human expert to ensure no offensive content is included in the DisastIR. We recognize potential risks associated with the misuse of retrieval models in disaster contexts, such as the spread of disinformation during crises. To mitigate these risks, DisastIR is intended solely for evaluation purposes and is released for research use only.
Acknowledgments
Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing
Appendix A Structural PDF File Processing Pipeline
Figure 6: Cohen’s kappa scores between LLM-based and human-annotated relevance labels across all LVHL datasets, as described in Section 4.2
All disaster management-related data (in PDF format) is obtained from publicly available sources with no personally identifiable information. Hence, explicit consent was not required. PDF files are collected using googlesearch-python (v1.3.0) and processed with PyMuPDF (v1.24.10) for content extraction. The extracted PDFs are then processed into text chunks through the following steps:
(1) Exact-URL Deduplication.
The URL of each downloaded PDF is recorded, and duplicate documents are removed by identifying identical download links.
(2) Text Extraction and Preprocessing.
Each PDF file is converted into plain text, where tables and figures are removed following the work of (Wen et al., 2023).
(3) Locality-Sensitive Hashing (LSH) Deduplication.
After cleaning, we apply LSH-based near-duplicate detection to identify and remove documents with highly overlapping content.
(4) Semantic Chunking.
Cleaned documents are segmented into semantically coherent text chunks. Each chunk is constrained to fewer than 256 tokens to optimize retrievability while maintaining semantic integrity.
(5) Embedding-based Near Deduplication.
To further eliminate redundancy at the passage level, dense embeddings are computed for all chunks. An ANN index is built to retrieve the top-$k$ nearest chunks, and pairs with cosine similarity above 0.9 are removed.
Appendix B Prompt Templates for Query Generation
Prompts for query generation based on disaster management-related passage under different search intents for QA, QAdoc, Twitter, FC, NLI, STS are in Tables 4, 5, 6, 7, 8, and 9.
Appendix C Prompt Templates for Relevance Labeling
This section presents the prompt templates used for LLM-based relevance judgments across six search intents, employing three prompting strategies: Zero-shot Direct Scoring, Chain-of-Thought Decomposed Reasoning, and Multi-Dimensional Attribute Scoring. For QA, QAdoc, and Twitter tasks, we adapt templates from Thomas et al. (2024); Farzi and Dietz (2024), as shown in Table 10. Based on these templates, we design relevance prompts for FC, NLI, and STS tasks, shown in Tables 11, 12, and 13, respectively. For STS, we adopt only Zero-shot Direct Scoring, as our preliminary experiments show it yields higher agreement with human labels (Cohen’s kappa). The estimated cost of generating 9,600 user queries and labeling over 1.3 million query-passage using GPT-4o-mini API is about \$1,400.
Appendix D Additional Analyses of Labeled query-passage pairs
As shown in Table 14, in certain retrieval tasks, such as NLI_Bio, NLI_Geo, the number of query-passage pairs assigned the highest relevance score is smaller than the number of user queries. This indicates that some queries do not have any passage in their candidate pool that is judged as fully relevant. For each query, we prompt an LLM to generate a directly relevant passage based on the associated domain passage and include it in the labeling pool (Section 3.4). However, the labeling results in Table 15 show that not all generated passages are considered fully relevant. This suggests that, even when guided by task-specific prompts, LLMs may produce passages that only partially address the query or fail to capture its key intent.
Many recent works have tried to employ LLM to generate synthetic training data to improve the quality of retrievers (Wang et al., 2023; Rajapakse and de Rijke, 2023; Xu et al., 2024; Lee et al., 2024b). This finding underscores the importance of consistency filtering (Alberti et al., 2019) to improve retrieval models’ performance, as LLM will generate irrelevant pairs. This aligns with prior research highlighting the need for consistency filtering when leveraging LLM-generated data to train retrievers Dai et al. (2022); Xu et al. (2024); Lee et al. (2024b).
Appendix E LVHL Dataset Construction
We use the names of 301 specific disaster event types as queries to search for disaster management-related user queries within each selected open-source dataset listed in Table 16. For each dataset, we first filter queries by keyword matching and then prompt an LLM to further remove queries that are irrelevant to disaster management. From the remaining queries, we randomly select up to 400 queries per dataset. The corresponding passage and relevance score in each source dataset are also included. This process results in the final query-passage pairs along with the human-annotated relevance scores used in the LVHL dataset for evaluating the agreement of LLM-based and human-annotated relevance scores.  [^4] LVHL is used solely to evaluate agreement between LLM and human annotations. It is not suitable for benchmarking retrieval models in the disaster management area, as most queries are drawn from training sets of the source datasets.
Appendix F LVHQ Dataset Construction
We sample 48 domain passages developed in Section 3.3, ensuring one passage per retrieval task and keeping all sampled passages different from those used in developing DisastIR. For each passage, a domain expert in the disaster management field is asked to read the passage and write a realistic user query that reflects a practical information need based on the content, resulting in 48 human-authored queries. The Human expert is given the same instructions for the query written (shown in Tables 4, 5, 6, 7, 8, 9) as those given to LLM to ensure fair comparison. In parallel, for the same set of passages, we also generate 48 queries using LLM in the same way as described in Section 3.4.
Each query, both human-authored and LLM-generated, is used to retrieve relevant passages from DisastIR corpus. As we have validated the agreement of LLM-based and Human-annotated relevance score in Section 4.2, all query-passage pairs are labeled in the same way as described in Section 3.5 and Section 3.6.
Figure 7: Performance of three top models across different tasks
Figure 8: Distribution of outliers of evaluated models’ performances
Appendix G Information of Evaluated Models and Model Implementation
Detailed information on all selected models is summarized in Table 17. The HuggingFace links and licenses of these models are in Table 18. The model parameter size is categorized as four levels: small (<109M), medium (109M - 305M), large (305M- 1B), and extra large (XL) (> 1B).
For each model, we follow official implementation guidelines to generate normalized query and passage embeddings. All evaluations are conducted in a zero-shot setting, with input sequences truncated to 512 tokens and a task-specific instruction prepended to each query. All models are run on a single NVIDIA A6000 GPU using HuggingFace Transformers, following the configurations specified in the official implementations.
Appendix H Performance of Evaluated Models
Performance of all evaluated models in all 48 search tasks in DisastIR is shown in Tables 19, 20, 21, 22, and 23.
Appendix I Additional Analyses of Model Performance across 48 Tasks
NV-Embed-v2 achieves the best performance on all QAdoc and NLI-related tasks (See Table 3 and Figure 5 in the main content). However, as shown in Figure 7, its poor results on Twitter-related tasks significantly lower its overall performance in DisastIR. This reflects its limitation in handling the informal, noisy, and contextually ambiguous nature of social media content. Given the importance of Twitter as a real-time, crowd-sourced information source during disasters (Alam et al., 2021; Yin et al., 2024; Lei et al., 2025), this weakness raises concerns about its reliability in real-world disaster response scenarios.
All three models perform poorly on NLI-related tasks, with the best achieving only an average score of 58.73 (Figure 7). Further analysis of outliers in the box plot (See Figure 3 in the main content) reveals that tasks causing significant performance drops consistently involve NLI search intents (Figure 8). This reveals a key limitation of current open-source SOTA retrievers, that they struggle with the complex reasoning required for NLI tasks in disaster contexts. Such limitations may lead to incorrect results or failure to retrieve critical information, which can negatively impact decision-making in disaster situations.
Information Need Generation Stage
"Brainstorm a list of useful text retrieval tasks where the goal is: Given a user question, retrieve passages that directly answer the question. Here are a few examples: Given a question about evacuation procedures during a flood, retrieve a passage that explains the steps involved. Given a question about the cause of infrastructure failure in a disaster, retrieve a passage identifying the cause. Given a question about relief funding timelines, retrieve a passage providing the relevant information. Guidelines: Each task description should be one sentence that clearly describes the user question and the kind of answer passages to be retrieved. Focus on real-world domains like disaster planning, relief logistics, early warning systems, community impact, government response, etc. Your output should be a JSON list of 3 strings, each describing a distinct and useful text retrieval task. Only output the list. Be creative. No explanations or additional content."
User Query Generation Stage
"You have been assigned a retrieval task: {task}. Your mission is to write one text retrieval example for this task in JSON format. The JSON object must contain the following keys:
  - user_query: a string, a random user search query specified by the retrieval task.
  - positive_document: a string, a relevant document for the user query.
Please adhere to the following guidelines:
  - The user_query should be {query_length}, {clarity}.
  - All documents must be created independent of the query. Avoid copying the query verbatim. It is acceptable if some parts of the positive_document are not topically related to the query.
  - All documents should be {num_words} long.
  - Do not provide any explanation in any document on why it is relevant or not relevant to the query. The query and documents must be realistic and inspired by real-world content (e.g., disaster management). All generated content should be in English no matter the provided content language is.
  - Both the query and documents require {difficulty} level education to understand.
Your output must always be a JSON object only, do not explain yourself or output anything else. Be creative!"
Table 4: Prompt templates for user query generation in QA-related tasks. The clarity placeholder takes values: clear, understandable with some effort, and ambiguous. The difficulty placeholder includes: elementary school, high school, college, and PhD. For query_length, possible values are: less than 10 words, 5 to 20 words, less than 20 words, at least 50 words, and at least 150 words. The num_words placeholder takes values such as: at least 100 words, at least 200 words, at most 50 words, and 50 to 150 words.
Information Need Generation Stage
"Brainstorm a list of document retrieval tasks, where the goal is: Given a user query, retrieve documents that provide useful and relevant answers. Here are a few examples to get you started: Given a query about emergency evacuation procedures, retrieve a document that outlines the proper steps. Given a query asking how heatwaves affect public health, retrieve a document discussing the medical or environmental impacts. Given a query about funding for post-disaster recovery, retrieve a document describing the financial aid process. Given a query on how early warning systems reduce disaster risk, retrieve a document explaining their function and benefits. Guidelines: Each task should be a single sentence describing what the query is and what kind of document should be retrieved in response. Tasks should span a broad range of information needs, from facts to procedures to causal relationships. Focus on disaster management-related themes such as risk mitigation, emergency logistics, climate impact, institutional roles, and infrastructure damage. Your output should be a JSON list of 20 strings, each describing a distinct and useful text retrieval task. Only output the list. No explanations."
User Query Generation Stage
"You have been assigned a document retrieval task: {task}. Your mission is to write one example for this task in JSON format. The JSON object must include:
  - user_query: a single, well-formed, natural language query that clearly asks for information based on the assigned task.
Guidelines: The query can be answered by the content provided in the following given paragraph. The query should reflect a realistic information need in the disaster management domain. Avoid generic or overly broad questions—make the query specific and grounded in actual scenarios (e.g., logistics, policies, actions). Use language inspired by real-world usage, such as what a policymaker, journalist, or emergency planner might ask. Output only a single JSON object with a user_query field. No extra formatting, documents, or explanations. Be clear, informative, and realistic!"
Table 5: Prompt templates for the user query generation for QAdoc-related search task. The clarity placeholder takes values: clear, understandable with some effort, and ambiguous. The difficulty placeholder includes: elementary school, high school, college, and PhD. For query_length, possible values are: less than 10 words, 5 to 20 words, less than 20 words, at least 50 words, and at least 150 words. The num_words placeholder takes values such as: at least 100 words, at least 200 words, at most 50 words, and 50 to 150 words.
Information Need Generation Stage
"Brainstorm a list of entity retrieval tasks where the goal is to retrieve tweets that mention or provide relevant information about one or more entities (e.g., organizations, people, locations, events) found in the query. Here are a few examples to inspire your thinking: Given a query referencing “UNICEF,” retrieve tweets about their emergency relief efforts. Given a query about “Cyclone Mocha,” retrieve tweets reporting its impact or aftermath. Given a query that mentions “World Health Organization,” retrieve tweets that discuss their role in disaster health responses. Given a query with “Manila,” retrieve tweets about disaster conditions or relief actions in that location. Guidelines: Each task should be a single sentence describing a situation where an entity is referenced and tweets related to that entity should be retrieved. Focus on disaster management-related entities such as emergency response agencies, international organizations, locations, events, or key figures. Encourage diversity in topics: ground response, aid distribution, weather events, infrastructure failure, etc. Your output should be a JSON list of about 3 strings, each describing a different NER Twitter retrieval task. No explanations or extra formatting. Be concise, diverse, and realistic."
User Query Generation Stage
"You have been assigned an entity-tweet retrieval task: {task}. Your mission is to write one example for this task in JSON format. The JSON object must include:
  - query: a sentence that mentions one or more disaster management-related entities.
  - positive_tweet: a tweet that provides relevant and informative content about the mentioned entity or entities.
Guidelines: The query should clearly mention a recognizable entity tied to the disaster domain. The positive_tweet should be informal, observational, or emotional—realistic Twitter-style language providing relevant details about the entity. All content should be inspired by disaster management-related themes such as rescue missions, weather events, humanitarian aid, response coordination, etc. Your output must be a single JSON object only. No explanation, no formatting beyond JSON. Keep it realistic and natural in tone."
Table 6: Prompt templates for the user query generation for Twitter-related search task. The clarity placeholder takes values: clear, understandable with some effort, and ambiguous. The difficulty placeholder includes: elementary school, high school, college, and PhD. For query_length, possible values are: less than 10 words, 5 to 20 words, less than 20 words, at least 50 words, and at least 150 words. The num_words placeholder takes values such as: at least 100 words, at least 200 words, at most 50 words, and 50 to 150 words.
Information Need Generation Stage
"Brainstorm a list of fact-checking retrieval tasks where the goal is: Given a claim, retrieve documents that either support or refute the claim, while distinguishing them from topically similar documents that do not address the claim’s veracity. Here are a few examples to guide your ideas: Given a claim about the effectiveness of early warning systems during floods, retrieve documents that either support or refute the claim. Given a claim about the number of people displaced by a recent earthquake, retrieve evidence that verifies or challenges it. Given a claim about the government’s relief distribution timeline, retrieve text that affirms or contradicts the stated timeline. Given a claim about the relationship between climate change and disaster frequency, retrieve relevant supporting or refuting content. Guidelines: Each task should be one sentence and describe what the claim is about and what kind of evidence is needed (support or refute). Base the topics on real-world domains such as disaster management, humanitarian aid, policy, climate, health impacts, etc. The tasks should vary in specificity and format (e.g., statistical claim, causal claim, factual assertion). Your output should be a JSON list of about 3 strings, each string representing a distinct fact-checking retrieval task. Output only the list. No explanations. Be creative and diverse in topic!"
User Query Generation Stage
"You have been assigned a fact-checking retrieval task: {task}. Your mission is to write one fact-checking retrieval instance in JSON format. The JSON object must contain:
  - claim: a short, factual or semi-factual statement (assertion) related to the task.
  - positive_document: a paragraph that supports or refutes the claim.
Guidelines: The claim should be {query_length}, {clarity}. The positive document must clearly support or refute the claim—either is acceptable. The claim should be clear, concise, and specific—not overly vague or too broad. Use examples from realistic disaster management-related content: climate events, emergency response, humanitarian relief, damage estimates, etc. All positive documents should be {num_words} long. All generated content should be in English no matter the provided content language is. Both the claim and documents must be understandable with {difficulty} level education. Output only a single JSON object. No additional text. Be precise and creative!"
Table 7: Prompt templates for the user query generation for fact-checking related search task. The clarity placeholder takes the values: clear, understandable with some effort, and ambiguous. The difficulty placeholder includes: elementary school, high school, college, and PhD. The query_length placeholder accepts values such as: less than 10 words, 5 to 20 words, at least 10 words, at least 20 words, and at least 50 words. The num_words placeholder includes: at most 15 words, at most 50 words, 50 to 150 words, at most 100 words, and at least 100 words.
Information Need Generation Stage
"Brainstorm a list of Natural Language Inference (NLI) retrieval tasks. In these tasks, the objective is: Given a premise sentence from a paragraph (e.g., about disaster management), retrieve a hypothesis sentence that is logically entailed by the premise. Here are a few examples to inspire your creativity: Given a sentence describing a government emergency response, retrieve a hypothesis that reflects an outcome or implication of that action. Given a statement about climate-induced hazards, retrieve a hypothesis summarizing the likely impact. Given a factual description of infrastructure damage, retrieve a hypothesis about the services affected. Given a claim about disaster preparedness strategies, retrieve a hypothesis that is logically supported by it. Guidelines: Each task description should be one sentence and should clearly specify the type of premise and the nature of the entailed hypothesis. Tasks should be generalizable across topics but inspired by domains such as climate, crisis response, risk, logistics, etc. Be diverse in topic and formality: from news-like to academic to conversational. Your output should be a JSON list of about 3 strings, each describing a different NLI retrieval task. Output only the list of task descriptions, no explanations."
User Query Generation Stage
"You have been assigned an NLI retrieval task: {task}. Your mission is to write one example for this task in JSON format. The JSON object must include:
  - premise: a sentence drawn or inspired from a paragraph (e.g., about disaster management).
  - entailed_hypothesis: a sentence that must logically follow from the premise.
Guidelines: The premise should be {query_length}, {clarity}. Use realistic examples from domains like climate risk, emergency response, infrastructure, health, or logistics, etc. Ensure the entailed hypothesis is non-trivial and clearly follows from the premise. Avoid word-for-word overlap between the sentences unless necessary for clarity. entailed_hypothesis should be {num_words} long. All generated content should be in English no matter the provided content language is. All contents require {difficulty} level education to understand and should be diverse in terms of topic and length. Output a single JSON object only. Do not explain yourself or add anything else. Be creative and accurate!"
Table 8: Prompt templates for the user query generation for NLI-related search task. The clarity placeholder takes values: clear, understandable with some effort, and ambiguous. The difficulty placeholder includes: elementary school, high school, college, and PhD. The query_length placeholder accepts values such as: less than 10 words, 5 to 20 words, at least 20 words, at least 50 words, and at least 150 words. The num_words placeholder includes: less than 10 words, 5 to 20 words, at least 20 words, at least 50 words, and at most 50 words.
Information Need Generation Stage
"Brainstorm a list of similar sentence retrieval tasks where the goal is: Given a sentence, retrieve other sentences that express the same or very similar meaning (paraphrases or semantically equivalent expressions). Here are a few examples to inspire your ideas: Given a sentence describing the impact of a flood, retrieve other sentences that paraphrase or closely restate the same impact. Given a sentence about the steps taken during emergency evacuation, retrieve sentences that express the same process using different wording. Given a sentence about climate-related disasters increasing in frequency, retrieve other sentences conveying the same trend. Given a factual statement about relief distribution, retrieve sentences that express the same fact using alternate phrasing. Guidelines: Each task should be written in one sentence and describe the kind of sentence (source) and what type of similar sentences should be retrieved. Focus on disaster management-related themes such as risk, policy, action, climate, or aid—but vary topics for diversity. Your output should be a JSON list of about 3 strings, each one describing a distinct similar sentence retrieval task. Output the list only. No explanations. Be creative and precise."
User Query Generation Stage
"You have been assigned a similar sentence retrieval task: {task}. Your mission is to write one example for this task in JSON format. The JSON object must contain:
  - query: a single sentence that expresses a specific idea.
  - positive: a sentence that expresses the same meaning as the query (semantic equivalence or high similarity).
Guidelines: The query should be {query_length}, {clarity}. The query and positive should be semantically equivalent, possibly using different wording or structure. Avoid copy-paste or trivial rewordings—be realistic and diverse. Use examples inspired by real-world disaster management-related content: emergency protocols, environmental impact, infrastructure, humanitarian response, etc. All positive documents should be {num_words} long. All generated content should be in English no matter the provided content language is. Both the query and documents must be understandable with {difficulty} level education. Output only a single JSON object. No explanation. Make it high-quality and realistic!"
Table 9: Prompt templates for the user query generation for STS-related search task. The clarity placeholder takes values: clear, understandable with some effort, and ambiguous. The difficulty placeholder includes: elementary school, high school, college, and PhD. The query_length placeholder accepts values such as: less than 10 words, 5 to 20 words, at least 50 words, and at most 50 words. The num_words placeholder includes: less than 10 words, 5 to 20 words, at least 50 words, and at most 50 words.
[1] Zero-shot Direct Scoring
You are a search quality rater evaluating passage relevance based on detailed instructions and outputting JSON.
Given a query and a passage, provide a score (0–3): 0 = Irrelevant, 1 = Related, 2 = Highly relevant, 3 = Perfectly relevant.
Important: 1 if somewhat related but not completely, 2 if important info + extra, 3 if only refers to topic.
Query: {query}, Passage: {passage}. Consider intent, content match (M), trustworthiness (T), then decide final score (O).
Output MUST be JSON: {"final_score": <0-3>}
[2] Chain-of-Thought Decomposed Reasoning
PHASE 1 – Answer Presence Prediction:
Instruction: Given a passage and a query, predict whether the passage includes an answer to the query by producing either “Yes” or “No”. Query: {query}, Passage: {passage}, Answer:
Output your prediction as JSON: {"has_answer": "Yes"} or {"has_answer": "No"}
PHASE 2 – Fine-grained Criterion Scoring:
System prompt: Please assess how well the provided passage meets specific criteria in relation to the query. Use the following scoring scale (0–3) for evaluation: 0: Not relevant at all / No information provided. 1: Marginally relevant / Partially addresses the criterion. 2: Fairly relevant / Adequately addresses the criterion. 3: Highly relevant / Fully satisfies the criterion.
Your output MUST be JSON: {"criterion_score": <0-3>}
User prompt: Please rate how well the given passage meets the {criterion_name} criterion in relation to the query. The output should be a single score (0–3) indicating {criterion_definition}. Query: {Query}, Passage: {Passage}, Output JSON: {"criterion_score": <integer_score_0_to_3>}
Note: placeholders for {criterion_name} and {criterion_definition} are:
{Exactness: "how precisely the passage answers the query", Coverage: "proportion of content discussing the query"}.
PHASE 3 – Final Relevance Scoring:
System prompt when Phase 1 = "Yes": You are a search quality rater. Provide a final relevance score (2 or 3):
2 = Highly relevant, 3 = Perfectly relevant. Your output MUST be JSON: {"relevance_score": <2_or_3>}
System prompt when Phase 1 = "No": You are a search quality rater. Provide a final relevance score (0 or 1).
0 = Irrelevant, 1 = Related. Your output MUST be JSON: {"relevance_score": <0_or_1>}
User prompt when Phase 1 = "Yes": The passage is relevant. Rate how relevant (2 or 3). Query: {Query}, Passage: {Passage}, Output JSON: {"relevance_score": <2_or_3>}
User prompt when Phase 1 = "No": The passage is irrelevant. Rate how irrelevant (0 or 1).Query: {Query}, Passage: {Passage}, Output JSON: {"relevance_score": <0_or_1>}
[3] Multi-Dimensional Attribute Scoring
PHASE 1 – Sub-criterion Scoring:
System prompt: Please assess how well the provided passage meets specific criteria in relation to the query. Use the following scoring scale (0–3): 0 = Not relevant at all / No information provided. 1 = Marginally relevant / Partially addresses the criterion. 2 = Fairly relevant / Adequately addresses the criterion. 3 = Highly relevant / Fully satisfies the criterion.
Your output MUST be JSON: {"criterion_score": <0-3>}
User prompt: Please rate how well the given passage meets the {criterion_name} criterion in relation to the query. The output should be a single score (0–3) indicating {criterion_definition}.
Query: {query}, Passage: {passage}, Output JSON: {"criterion_score": <integer_score_0_to_3>}
Note: placeholders for {criterion_name} and {criterion_definition} are:
{Exactness: "how precisely the passage answers the query", Coverage: "proportion of content discussing the query", Topicality: "subject alignment between passage and query", Contextual Fit: "presence of relevant background information"}.
PHASE 2 – Final Relevance Aggregation:
System prompt: You are a search quality rater evaluating overall relevance. Given a query, passage, and sub-scores, provide a final score (0–3): 3 = Perfectly relevant, 2 = Highly relevant, 1 = Related, 0 = Irrelevant. Your output MUST be JSON: {"final_relevance_score": <0-3>}
User prompt: Please rate how relevant the passage is based on the given sub-scores.Query: {query}, Passage: {passage}, Exactness: LLM score, Coverage: LLM score, Topicality: LLM score, Contextual Fit: LLM score.
Output your final rating as JSON: {"final_relevance_score": <integer_score_0_to_3>}
Table 10: LLM relevance judgment prompt templates for QA, QAdoc, and Twitter-related search tasks.
[1] Zero-shot Direct Scoring
System prompt: You are a search quality rater evaluating evidence for fact-checking and outputting JSON.
User prompt: Given a claim and a passage, score relevance (003333): 00 = Irrelevant, $1$ = Related (no help), $2$ = Relevant (unclear/mixed), $3$ = Direct support/refutation.
Important: $1$ if related but no help, $2$ if important info + noise, $3$ if clearly supports/refutes.
Claim: {Query}, Passage: {Passage}. Consider intent, support/refutation ($M$), trustworthiness ($T$), then decide final score ($O$). Output MUST be JSON: {"final_score": <0_to_3>}
[2] Chain-of-Thought Decomposed Reasoning
PHASE 1 – Answer Presence Prediction:
Instruction: Given a passage and a claim, predict whether the passage includes information that supports or refutes the claim by producing either “Yes” or “No”.
Claim: {Query}, Passage: {Passage}. Output JSON: {"has_answer": "Yes"} or {"has_answer": "No"}
PHASE 2 – Fine-grained Criterion Scoring:
System prompt: Please assess how well the provided passage serves as evidence for evaluating the claim. Use the following scoring scale (003333): 00 = Not relevant at all / No information provided. $1$ = Marginally relevant / Partially addresses the criterion.2222 = Fairly relevant / Adequately addresses the criterion. $3$ = Highly relevant / Fully satisfies the criterion. Your output MUST be JSON: {"criterion_score": <0_to_3>}
User prompt: Please rate how well the given passage meets the {criterion_name} criterion in relation to the claim. The output should be a single score (003333) indicating {criterion_definition}.
Claim: {Query}, Passage: {Passage}. Output JSON: {"criterion_score": <integer_score_0_to_3>}
Note: placeholders for {criterion_name} and {criterion_definition} are:
{Exactness: "how precisely the passage supports or refutes the claim", Coverage: "the extent to which the passage discusses content directly relevant to the claim",. Each criterion evaluated independently.
PHASE 3 – Final Relevance Scoring:
System prompt when Phase 1 = "Yes": You are a rater evaluating evidence for fact-checking. Score ($2$ or $3$): $2$ = Highly relevant (some support/refutation), $3$ = Perfectly relevant (direct support/refutation). Your output MUST be JSON: {"relevance_score": <2_or_3>} System prompt when Phase 1 = "No": You are a rater evaluating evidence for fact-checking. Score (00 or $1$): 00 = Irrelevant, $1$ = Related (no support/refutation). Your output MUST be JSON: {"relevance_score": <0_or_1>}
User prompt when Phase 1 = "Yes": Passage is relevant. Rate its relevance ($2$ or $3$). Claim: {Query}, Passage: {Passage}. Output JSON: {"relevance_score": <2_or_3>}
User prompt when Phase 1 = "No":
Passage irrelevant for evidence. Rate its relevance (00 or $1$). Claim: {Query}, Passage: {Passage}. Output JSON: {"relevance_score": <0_or_1>}
[3] Multi-Dimensional Attribute Scoring
PHASE 1 – Sub-criterion Scoring:
System prompt: Please assess how well the provided passage serves as evidence for evaluating the claim according to the following criteria. Use a score (003333): 00 = Not relevant at all / No information provided. $1$ = Marginally relevant / Partially addresses the criterion. $2$ = Fairly relevant / Adequately addresses the criterion. $3$ = Highly relevant / Fully satisfies the criterion. Your output MUST be JSON: {"criterion_score": <0_to_3>}
User prompt: Please rate how well the given passage meets the {criterion_name} criterion in relation to the claim. The output should be a single score (003333) indicating {criterion_definition}.
Claim: {Query}, Passage: {Passage}. Output JSON: {"criterion_score": <0_to_3>}
Note: placeholders for {criterion_name} and {criterion_definition} are:
{Exactness: "how precisely the passage supports or refutes the claim", Coverage: "the extent to which the passage discusses content directly relevant to the claim", Topicality: "how closely the subject matter aligns with the claim topic", Contextual Fit: "how much relevant background/context is provided to verify the claim"}.
PHASE 2 – Final Relevance Aggregation:
System prompt: You are a search quality rater evaluating evidence relevance for fact-checking. Given a claim, passage, and sub-scores, provide a final score (003333): $3$ = Perfectly relevant (direct support/refutation), $2$ = Highly relevant (helps verification), $1$ = Related (topic match, no verification aid), 00 = Irrelevant. Your output MUST be JSON: {"final_relevance_score": <0_to_3>}
User prompt: Please rate how relevant the given passage is to the claim based on the given scores.
Claim: {Query}, Passage: {Passage}, Exactness: LLM score, Coverage: LLM score, Topicality: LLM score, Contextual Fit: LLM score. Output JSON: {"final_relevance_score": <0_to_3>}
Table 11: LLM relevance judgment prompt templates for FC-related search tasks.
[1] Zero-shot Direct Scoring
System prompt: You are a rater evaluating entailment. Output only the final score in JSON.
User prompt: Given a premise and hypothesis, score entailment (003333): 00 = Not entailed/Contradicted, $1$ = Related not entailed, $2$ = Mostly entailed, $3$ = Perfectly entailed.
Important: $1$ if related but not inferable, $2$ if captures important implied content but not fully, $3$ if clearly/fully supported.
Premise: {Query}, Hypothesis: {Passage}. Consider premise implications, logical following ($E$), info gaps, then decide final score ($O$). Output MUST be JSON: {"final_score": <0_to_3>}
[2] Chain-of-Thought Decomposed Reasoning
PHASE 1 – Answer Presence Prediction:
Instruction: Given a hypothesis and a premise, predict whether the hypothesis is entailed by the premise by producing either “Yes” or “No”.
Premise: {Query}, Hypothesis: {Passage}.Output JSON: {"has_answer": "Yes"} or {"has_answer": "No"}
PHASE 2 – Fine-grained Criterion Scoring:
System prompt: Please assess how well the given hypothesis meets the {criterion_name} criterion in relation to the premise.
Use a single score (003333) indicating {criterion_definition}.
Premise: {Query}, Hypothesis: {Passage}. Output JSON: {"criterion_score": <0_to_3>}
Note: placeholders for {criterion_name} and {criterion_definition} are: {Exactness: "how precisely the hypothesis is entailed by the premise", Coverage: "the extent to which the hypothesis reflects core information from the premise".
PHASE 3 – Final Relevance Scoring:
System prompt when Phase 1 = "Yes": You are a rater evaluating entailment. Provide a final score ($2$ or $3$): $2$ = Mostly entailed, $3$ = Perfectly entailed. Your output MUST be JSON: {"relevance_score": <2_or_3>}
System prompt when Phase 1 = "No": You are a rater evaluating entailment. Provide a final score (00 or $1$): 00 = Not entailed/Contradicted, $1$ = Related but not entailed. Your output MUST be JSON: {"relevance_score": <0_or_1>}
User prompt when Phase 1 = "Yes": Hypothesis is entailed. Rate how well ($2$ or $3$). Premise: {Query}, Hypothesis: {Passage}. Output JSON: {"relevance_score": <2_or_3>}
User prompt when Phase 1 = "No": Hypothesis is not entailed. Rate how (00 or $1$). Premise: {Query}, Hypothesis: {Passage}. Output JSON: {"relevance_score": <0_or_1>}
[3] Multi-Dimensional Attribute Scoring
PHASE 1 – Sub-criterion Scoring:
System prompt: Please assess how well the hypothesis is entailed by the premise according to the following criteria. Use a score (003333): 00 = Not relevant at all / No information provided. $1$ = Marginally relevant / Partially addresses the criterion. $2$ = Fairly relevant / Adequately addresses the criterion. $3$ = Highly relevant / Fully satisfies the criterion. Your output MUST be JSON: {"criterion_score": <0_to_3>}
User prompt: Please rate how well the given hypothesis meets the {criterion_name} criterion in relation to the premise. The output should be a single score (003333) indicating {criterion_definition}.
Premise: {Query}, Hypothesis: {Passage}. Output JSON: {"criterion_score": <0_to_3>}
Note: placeholders for {criterion_name} and {criterion_definition} are: {Exactness: "how precisely the hypothesis is entailed by the premise", Coverage: "the extent to which the hypothesis reflects core information from the premise", Topicality: "how closely the subject matter of the hypothesis aligns with that of the premise", Contextual Fit: "how well the hypothesis fits within the context or background established by the premise"}.
PHASE 2 – Final Relevance Aggregation:
System prompt: You are a search quality rater evaluating entailment. Given a premise, hypothesis, and sub-scores, provide a final score (003333): $3$ = Perfectly entailed, $2$ = Mostly entailed, $1$ = Related but not entailed, 00 = Not entailed/Contradicted. Your output MUST be JSON: {"final_relevance_score": <0_to_3>}
User prompt: Please rate how well the hypothesis is entailed by the premise based on the given scores.
Premise: {Query}, Hypothesis: {Passage}, Exactness: LLM score, Coverage: LLM score, Topicality: LLM score, Contextual Fit: LLM score. Output JSON: {"final_relevance_score": <0_to_3>}
Table 12: LLM relevance judgment prompt templates for NLI-related search tasks.
[1] Zero-shot Direct Scoring for STS
System prompt: You are a semantic-similarity rater. Output JSON with a score (005555) where 00 = unrelated, $5$ = semantically equivalent.
User prompt: Rate the semantic similarity 005555: 00 = unrelated $\mid$ $1$ = slight $\mid$ $2$ = partial $\mid$ $3$ = moderate $\mid$ $4$ = high $\mid$ $5$ = equivalent
Sentence A: {input1} Sentence B: {input2}
Output JSON: {"final_similarity_score": <0_to_5>}
Table 13: LLM relevance judgment prompt templates for STS-related search tasks.
Search Intent Event Type rel=0 rel=1 rel=2 rel=3 rel=4 rel=5
FC Bio 17987 3034 1952 1014 0 0
Chem 16345 3845 2746 1656 0 0
Env 16739 4243 3235 1588 0 0
Extra 16867 3748 2436 1312 0 0
Geo 20151 4126 2549 1038 0 0
MH 18671 4283 2601 1115 0 0
Soc 20757 3287 2707 1099 0 0
Tech 18651 3905 2958 1245 0 0
NLI Bio 19031 5388 1284 193 0 0
Chem 19391 5981 2151 333 0 0
Env 16130 6823 2041 213 0 0
Extra 19372 5437 1378 212 0 0
Geo 20557 6098 1404 151 0 0
MH 19151 6100 1587 214 0 0
Soc 19597 5570 1612 218 0 0
Tech 21240 5540 1430 184 0 0
QA Bio 19843 3073 3099 636 0 0
Chem 18823 3411 3662 989 0 0
Env 18073 3838 4036 738 0 0
Extra 19293 3480 3300 734 0 0
Geo 19478 3699 3373 590 0 0
MH 20241 3819 3689 673 0 0
Soc 19832 3206 3551 527 0 0
Tech 19803 4005 3731 505 0 0
QAdoc Bio 18615 4424 2112 184 0 0
Chem 18254 4871 2604 303 0 0
Env 16710 5900 3125 195 0 0
Extra 17120 5394 2749 335 0 0
Geo 19182 5157 2075 159 0 0
MH 19610 4898 2537 211 0 0
Soc 15153 4981 2937 282 0 0
Tech 19247 5042 2569 213 0 0
STS Bio 9189 8880 5229 2303 1350 114
Chem 8201 8400 5780 2670 1605 131
Env 7276 8600 6381 2938 1744 109
Extra 9625 8639 5151 2459 1322 117
Geo 8815 10232 6709 2771 1182 107
MH 8590 8809 6436 3072 1680 115
Soc 9402 8349 5114 2575 1520 114
Tech 7846 9316 5606 2686 1345 121
Twitter Bio 29385 3650 1904 243 0 0
Chem 27326 4259 2312 289 0 0
Env 24757 5062 3058 366 0 0
Extra 26604 4168 2084 346 0 0
Geo 28491 4482 2257 273 0 0
MH 27086 4255 2252 331 0 0
Soc 28238 3472 1861 263 0 0
Tech 26349 4457 2219 363 0 0
Table 14: Distribution of qrels scores rel=0 through rel=5 for each search task in DisastIR. “rel” represents relevance score. Only STS-related search task is labeled in 6 levels, with others labeled in 4 levels.
Search Intent Event Type rel=0 rel=1 rel=2 rel=3 rel=4 rel=5
FC Bio 1 11 19 169 0 0
Chem 0 5 15 180 0 0
Env 0 2 14 184 0 0
Extra 1 6 24 169 0 0
Geo 1 9 21 169 0 0
MH 0 4 26 170 0 0
Soc 1 2 17 180 0 0
Tech 0 6 16 178 0 0
NLI Bio 8 33 78 81 0 0
Chem 7 39 81 73 0 0
Env 10 29 96 65 0 0
Extra 20 40 79 61 0 0
Geo 6 51 89 54 0 0
MH 10 44 67 79 0 0
Soc 5 52 77 66 0 0
Tech 13 39 84 64 0 0
QA Bio 1 3 33 163 0 0
Chem 0 3 30 167 0 0
Env 0 2 36 162 0 0
Extra 0 4 27 169 0 0
Geo 2 1 38 159 0 0
MH 0 2 38 160 0 0
Soc 0 5 30 165 0 0
Tech 0 2 48 150 0 0
QAdoc Bio 3 38 96 63 0 0
Chem 6 45 85 64 0 0
Env 1 35 102 62 0 0
Extra 4 34 77 85 0 0
Geo 2 44 81 73 0 0
MH 4 36 74 86 0 0
Soc 2 19 95 84 0 0
Tech 3 28 79 90 0 0
STS Bio 0 0 4 8 79 109
Chem 0 0 1 3 80 116
Env 0 0 1 7 88 104
Extra 0 0 0 5 85 110
Geo 0 0 0 4 94 102
MH 0 0 0 6 88 106
Soc 0 0 0 11 87 102
Tech 0 0 2 8 83 107
Twitter Bio 0 10 94 96 0 0
Chem 1 19 98 82 0 0
Env 0 12 103 85 0 0
Extra 0 14 101 85 0 0
Geo 0 9 95 96 0 0
MH 4 9 74 113 0 0
Soc 0 16 104 80 0 0
Tech 0 19 94 87 0 0
Table 15: Distribution of query-generated relevant document relevance scores rel-0 through rel-5
Name Intent Link #
MS Macro QA https://huggingface.co/datasets/microsoft/ms_marco 400
TriviaQA QA https://huggingface.co/datasets/sentence-transformers/trivia-qa-triplet 400
ALLNLI NLI https://huggingface.co/datasets/sentence-transformers/all-nli 400
XNLI NLI https://huggingface.co/datasets/mteb/xnli/viewer/en 400
STSB STS https://huggingface.co/datasets/sentence-transformers/stsb 400
ClimateFever FC https://huggingface.co/datasets/tdiggelm/climate_fever 400
Table 16: Overview of selected open-source datasets in LVHL. “#” represents the number of selected queries in the corresponding dataset.
Model Param Size Base Embed. MTEB Arch.
Name Size Bin Model Size Rank
inf-retriever-v1 7B XL gte-Qwen2-7B-instruct 3584 1 decoder
NV-Embed-v2 7B XL Mistral-7B-v0.1 4096 2 decoder
inf-retriever-v1-1.5b 1.5B XL gte-Qwen2-1.5B-instruct 1536 3 decoder
Linq-Embed-Mistral 7B XL E5-mistral-7b-instruct 4096 4 decoder
NV-Embed-v1 7B XL Mistral-7B-v0.1 4096 5 decoder
SFR-Embedding-Mistral 7B XL E5-mistral-7b-instruct 4096 6 decoder
snowflake-arctic-embed-l 335M Large e5-large-unsupervised 1024 7 encoder
snowflake-arctic-embed-l-v2.0 568M Large gte-multilingual-mlm-base 1024 8 encoder
snowflake-arctic-embed-m-v2.0 305M Medium bge-m3-retromae 768 9 encoder
gte-Qwen2-7B-instruct 7B XL Qwen2-7B 3584 10 decoder
snowflake-arctic-embed-m-v1.5 109M Medium BERT-base-uncased 768 11 encoder
e5-mistral-7b-instruct 7B XL Mistral-7b 4096 12 decoder
snowflake-arctic-embed-m 109M Medium e5-unsupervised-base 764 13 encoder
granite-embedding-125m-english 125M Medium RoBERTa 768 14 encoder
bge-large-en-v1.5 335M Large 1024 15 encoder
mxbai-embed-large-v1 335M Large 1024 16 encoder
snowflake-arctic-embed-s 33M Small e5-unsupervised-small 384 17 encoder
bge-base-en-v1.5 109M Medium 768 18 encoder
bge-small-en-v1.5 33M Small 384 19 encoder
multilingual-e5-large-instruct 560M Large xlm-roberta-large 1024 20 encoder
thenlper-gte-base 109M Medium EBRT-base 768 21 encoder
multilingual-e5-large 560M Large xlm-roberta-large 1024 22 encoder
gte-Qwen2-1.5B-instruct 1.5B XL Qwen2-1.5B 1536 23 decoder
e5-base-v2 109M Medium bert-large-uncased 1024 24 encoder
e5-large-v2 335M Large bert-base-uncased 768 25 encoder
e5-small-v2 33M Small MiniLM 384 26 encoder
all-mpnet-base-v2 109M Medium microsoft/mpnet-base 768 27 encoder
gte-base-en-v1.5 137M Medium EBRT-base 768 28 encoder
gte-large-en-v1.5 434M Large EBRT-large 1024 encoder
llmrails-ember-v1 335M Large 1024 encoder
Table 17: Information of all evaluated models. “–” means no publicly available information is available.
Model Name Link License
inf-retriever-v1 https://huggingface.co/infly/inf-retriever-v1 apache-2.0
NV-Embed-v2 https://huggingface.co/nvidia/NV-Embed-v2 cc-by-nc-4.0
inf-retriever-v1-1.5b https://huggingface.co/infly/inf-retriever-v1-1.5b apache-2.0
Linq-Embed-Mistral https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral cc-by-nc-4.0
NV-Embed-v1 https://huggingface.co/nvidia/NV-Embed-v1 cc-by-nc-4.0
SFR-Embedding-Mistral https://huggingface.co/Salesforce/SFR-Embedding-Mistral cc-by-nc-4.0
snowflake-arctic-embed-l https://huggingface.co/Snowflake/snowflake-arctic-embed-l apache-2.0
snowflake-arctic-embed-l-v2.0 https://huggingface.co/Snowflake/snowflake-arctic-embed-l-v2.0 apache-2.0
snowflake-arctic-embed-m-v2.0 https://huggingface.co/Snowflake/snowflake-arctic-embed-m-v2.0 apache-2.0
gte-Qwen2-7B-instruct https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct apache-2.0
snowflake-arctic-embed-m-v1.5 https://huggingface.co/Snowflake/snowflake-arctic-embed-m-v1.5 apache-2.0
e5-mistral-7b-instruct https://huggingface.co/intfloat/e5-mistral-7b-instruct mit
snowflake-arctic-embed-m https://huggingface.co/Snowflake/snowflake-arctic-embed-m apache-2.0
granite-embedding-125m-english https://huggingface.co/ibm-granite/granite-embedding-125m-english mit
bge-large-en-v1.5 https://huggingface.co/BAAI/bge-large-en-v1.5 apache-2.0
mxbai-embed-large-v1 https://huggingface.co/mixedbread-ai/mxbai-embed-large-v1 apache-2.0
snowflake-arctic-embed-s https://huggingface.co/Snowflake/snowflake-arctic-embed-s mit
bge-base-en-v1.5 https://huggingface.co/BAAI/bge-base-en-v1.5 mit
bge-small-en-v1.5 https://huggingface.co/BAAI/bge-small-en-v1.5 mit
multilingual-e5-large-instruct https://huggingface.co/intfloat/multilingual-e5-large-instruct mit
thenlper-gte-base https://huggingface.co/thenlper/gte-base mit
multilingual-e5-large https://huggingface.co/intfloat/multilingual-e5-large-instructt apache-2.0
gte-Qwen2-1.5B-instruct https://huggingface.co/Alibaba-NLP/gte-Qwen2-1.5B-instruct mit
e5-base-v2 https://huggingface.co/intfloat/e5-base-v2 mit
e5-large-v2 https://huggingface.co/intfloat/e5-large-v2 mit
e5-small-v2 https://huggingface.co/intfloat/e5-small-v2 apache-2.0
all-mpnet-base-v2 https://huggingface.co/sentence-transformers/all-mpnet-base-v2 apache-2.0
gte-base-en-v1.5 https://huggingface.co/Alibaba-NLP/gte-base-en-v1.5 apache-2.0
gte-large-en-v1.5 https://huggingface.co/Alibaba-NLP/gte-large-en-v1.5 mit
llmrails-ember-v1 https://huggingface.co/llmrails/ember-v1 mit
Table 18: HuggingFace model links and licenses for all evaluated models.
QA QAdoc Twitter FC NLI STS Avg.
Alibaba-NLP-gte-Qwen2-1.5B-instruct
Biological 12.93 19.75 19.16 25.18 18.82 34.19 21.67
Chemical 14.74 20.20 19.57 25.92 15.82 26.73 20.50
Environmental 14.03 26.03 23.66 21.48 19.75 34.39 23.23
Extraterrestrial 13.13 21.89 15.62 27.12 17.38 30.09 20.87
Geohazard 14.68 17.35 18.51 21.79 15.51 34.44 20.38
Meteorological&hydrological 12.63 19.33 23.63 22.10 19.47 30.22 21.23
Societal 14.51 29.00 17.56 24.04 18.59 22.57 21.05
Technological 15.20 24.16 19.15 23.58 18.63 37.00 22.95
Avg. 13.98 22.21 19.61 23.90 18.00 31.20 21.48
Alibaba-NLP-gte-Qwen2-7B-instruct
Biological 62.58 37.35 43.67 27.64 45.76 71.38 48.06
Chemical 64.42 38.19 41.06 31.31 46.84 70.43 48.71
Environmental 62.58 38.36 41.83 29.77 48.23 71.06 48.64
Extraterrestrial 58.55 34.42 37.55 31.45 46.54 73.40 46.99
Geohazard 61.12 36.31 40.13 28.31 47.12 68.06 46.84
Meteorological&hydrological 62.22 39.15 42.08 26.14 45.62 70.57 47.63
Societal 66.77 47.63 43.10 31.95 51.06 72.05 52.09
Technological 61.51 41.51 41.80 33.27 47.08 70.86 49.34
Avg. 62.47 39.12 41.40 29.98 47.28 70.98 48.54
Alibaba-NLP-gte-base-en-v1.5
Biological 60.86 54.84 46.25 53.77 42.53 70.95 54.87
Chemical 60.99 55.23 43.96 53.63 37.30 70.23 53.56
Environmental 62.02 55.77 47.11 51.32 43.30 71.72 55.21
Extraterrestrial 59.48 53.38 49.06 53.30 37.65 72.35 54.20
Geohazard 60.10 54.26 45.96 49.82 36.83 70.45 52.90
Meteorological&hydrological 59.98 58.31 47.26 51.44 39.13 70.93 54.51
Societal 61.99 59.05 48.04 53.11 42.88 71.83 56.15
Technological 59.83 57.66 48.85 53.55 39.39 71.68 55.16
Avg. 60.66 56.06 47.06 52.49 39.88 71.27 54.57
Alibaba-NLP-gte-large-en-v1.5
Biological 68.22 59.69 43.28 56.01 36.94 67.79 55.32
Chemical 70.28 59.30 38.93 56.44 33.85 68.50 54.55
Environmental 67.62 57.23 41.34 52.00 38.74 66.61 53.92
Extraterrestrial 65.28 56.57 42.45 51.24 32.11 68.85 52.75
Geohazard 65.65 52.61 35.16 49.45 30.81 64.75 49.74
Meteorological&hydrological 67.20 59.98 38.81 51.14 33.15 66.41 52.78
Societal 69.23 62.26 41.72 53.73 38.01 68.27 55.54
Technological 68.15 60.74 40.77 54.70 34.90 67.23 54.41
Avg. 67.70 58.55 40.31 53.09 34.81 67.30 53.63
BAAI-bge-base-en-v1.5
Biological 51.58 53.85 50.86 58.41 37.78 65.45 52.99
Chemical 47.97 50.77 48.75 58.64 34.71 67.10 51.32
Environmental 50.63 51.44 47.75 60.93 41.00 63.61 52.56
Extraterrestrial 50.50 57.93 45.21 59.49 33.18 65.59 51.98
Geohazard 47.81 52.76 44.28 56.91 33.09 62.52 49.56
Meteorological&hydrological 49.27 53.06 47.10 59.05 34.13 62.09 50.78
Societal 50.00 59.17 45.32 55.97 36.82 62.81 51.68
Technological 49.56 52.77 46.99 58.39 33.90 66.24 51.31
Avg. 49.67 53.97 47.03 58.47 35.58 64.43 51.52
BAAI-bge-large-en-v1.5
Biological 56.35 55.10 36.71 55.50 35.62 66.83 51.02
Chemical 56.24 51.49 33.57 56.73 34.65 66.02 49.78
Environmental 57.20 51.53 33.76 54.54 38.85 65.74 50.27
Extraterrestrial 58.77 57.02 33.03 58.30 32.68 66.38 51.03
Geohazard 55.15 53.62 30.65 53.17 35.16 63.51 48.54
Meteorological&hydrological 56.45 54.24 31.36 53.73 33.20 62.50 48.58
Societal 58.61 58.52 32.33 54.25 35.96 63.75 50.57
Technological 57.39 56.25 30.28 55.56 35.44 66.37 50.21
Avg. 57.02 54.72 32.71 55.22 35.19 65.14 50.00
Table 19: Performance of the first six evaluated models under six search intents and eight event types under the exact search setting. Part I
QA QAdoc Twitter FC NLI STS Avg.
BAAI-bge-small-en-v1.5
Biological 47.81 47.68 31.38 53.01 28.93 60.77 44.93
Chemical 45.25 42.32 30.68 55.45 27.44 57.61 43.12
Environmental 51.27 45.90 28.36 52.48 31.09 53.90 43.83
Extraterrestrial 48.32 45.50 26.64 53.35 25.57 58.43 42.97
Geohazard 46.74 42.91 25.79 50.88 25.91 55.30 41.25
Meteorological&hydrological 46.89 46.96 26.52 52.61 28.11 55.46 42.76
Societal 47.90 47.18 26.05 45.50 25.78 54.76 41.20
Technological 44.82 45.61 27.12 53.74 24.75 58.68 42.45
Avg. 47.38 45.51 27.82 52.13 27.20 56.86 42.82
Linq-AI-Research-Linq-Embed-Mistral
Biological 35.16 31.88 27.70 43.16 28.81 50.09 36.13
Chemical 39.05 28.91 27.52 44.70 20.32 39.50 33.33
Environmental 42.55 33.30 30.97 36.88 30.97 50.49 37.53
Extraterrestrial 33.98 33.26 25.64 44.49 28.29 42.68 34.72
Geohazard 35.77 28.05 23.63 33.08 27.89 47.75 32.70
Meteorological&hydrological 33.17 28.16 28.04 32.61 28.43 44.46 32.48
Societal 35.09 36.96 23.60 38.97 30.61 39.84 34.18
Technological 30.12 30.78 25.34 41.78 23.06 50.72 33.63
Avg. 35.61 31.41 26.55 39.46 27.30 45.69 34.34
Salesforce-SFR-Embedding-Mistral
Biological 72.89 67.94 72.41 69.18 51.66 74.32 68.07
Chemical 72.03 66.49 69.62 72.05 49.51 72.65 67.06
Environmental 72.25 66.24 72.43 73.16 52.99 73.20 68.38
Extraterrestrial 70.01 68.71 67.85 70.71 47.73 75.48 66.75
Geohazard 70.77 65.41 69.06 69.21 49.01 70.81 65.71
Meteorological&hydrological 72.44 67.29 70.11 70.49 50.84 73.12 67.38
Societal 72.78 70.13 71.38 69.38 53.92 74.26 68.64
Technological 70.89 68.42 70.48 70.80 51.20 75.08 67.81
Avg. 71.76 67.58 70.42 70.62 50.86 73.61 67.47
Snowflake-snowflake-arctic-embed-l
Biological 41.34 30.98 16.73 33.28 35.21 54.92 35.41
Chemical 40.77 30.26 16.39 34.55 33.24 56.39 35.27
Environmental 43.56 34.71 17.68 34.42 40.28 56.81 37.91
Extraterrestrial 41.09 27.96 14.48 31.93 32.31 58.93 34.45
Geohazard 36.34 28.20 12.92 33.83 31.08 53.90 32.71
Meteorological&hydrological 37.75 25.59 14.99 33.70 34.90 57.75 34.11
Societal 42.99 35.02 16.49 28.39 36.37 57.69 36.16
Technological 42.70 30.53 12.85 31.51 33.57 58.20 34.89
Avg. 40.82 30.41 15.32 32.70 34.62 56.82 35.11
Snowflake-snowflake-arctic-embed-l-v2.0
Biological 53.33 58.87 38.42 60.63 41.22 62.01 52.41
Chemical 56.40 57.57 40.62 61.11 39.01 63.87 53.10
Environmental 57.52 59.91 45.04 62.53 45.87 63.73 55.77
Extraterrestrial 54.74 56.70 36.38 60.14 37.86 62.82 51.44
Geohazard 54.21 57.26 34.76 59.82 39.80 60.19 51.01
Meteorological&hydrological 51.69 59.76 36.28 61.42 40.67 62.49 52.05
Societal 57.78 63.46 41.87 57.77 44.56 64.80 55.04
Technological 57.02 61.94 36.67 59.97 40.49 67.14 53.87
Avg. 55.34 59.43 38.75 60.42 41.19 63.38 53.09
Snowflake-snowflake-arctic-embed-m
Biological 31.68 13.54 8.09 38.04 39.39 54.55 30.88
Chemical 33.35 14.43 9.76 35.38 36.26 58.17 31.23
Environmental 35.53 14.87 9.18 35.83 43.61 57.19 32.70
Extraterrestrial 33.40 14.31 8.29 35.18 39.67 58.59 31.57
Geohazard 31.14 14.32 6.15 35.78 36.75 54.06 29.70
Meteorological&hydrological 30.96 14.49 9.29 35.75 39.04 56.34 30.98
Societal 35.17 14.34 10.72 31.87 40.57 57.82 31.75
Technological 35.55 13.70 7.63 34.58 36.14 58.32 30.99
Avg. 33.35 14.25 8.64 35.30 38.93 56.88 31.23
Table 20: Performance of evaluated models under six search intents and eight event types under the exact search setting. Part II
QA QAdoc Twitter FC NLI STS Avg.
Snowflake-snowflake-arctic-embed-m-v1.5
Biological 20.41 30.10 17.56 48.74 42.93 63.69 37.24
Chemical 23.39 29.15 20.20 48.55 41.74 65.45 38.08
Environmental 28.64 33.80 21.74 51.30 49.00 65.53 41.67
Extraterrestrial 25.43 24.07 17.35 46.65 41.79 66.04 36.89
Geohazard 26.06 30.31 17.01 46.63 42.02 63.49 37.59
Meteorological&hydrological 22.61 28.98 17.43 48.37 42.51 64.42 37.39
Societal 29.55 35.48 17.74 47.59 42.61 64.06 39.51
Technological 29.77 32.56 17.47 48.12 41.71 66.91 39.42
Avg. 25.73 30.56 18.31 48.24 43.04 64.95 38.47
Snowflake-snowflake-arctic-embed-m-v2.0
Biological 60.58 62.53 48.09 57.58 42.68 64.84 56.05
Chemical 61.22 61.01 47.12 59.78 40.78 64.72 55.77
Environmental 63.96 62.03 49.99 60.78 47.75 65.53 58.34
Extraterrestrial 60.83 61.17 46.86 58.68 39.05 65.74 55.39
Geohazard 60.38 60.30 47.90 57.05 43.33 63.02 55.33
Meteorological&hydrological 60.47 63.37 46.95 58.29 40.69 63.79 55.59
Societal 62.72 65.79 47.63 54.61 43.17 66.73 56.78
Technological 61.34 63.79 47.86 57.56 41.24 67.88 56.61
Avg. 61.44 62.50 47.80 58.04 42.34 65.28 56.23
Snowflake-snowflake-arctic-embed-s
Biological 34.80 26.88 18.93 44.11 36.25 66.25 37.87
Chemical 33.78 28.03 20.01 45.40 35.16 66.72 38.18
Environmental 39.09 30.43 23.33 42.75 43.67 68.17 41.24
Extraterrestrial 35.15 21.60 16.03 41.94 34.80 66.06 35.93
Geohazard 35.78 25.98 16.63 41.53 36.37 63.31 36.60
Meteorological&hydrological 34.13 28.33 16.74 43.00 38.08 66.33 37.77
Societal 37.91 30.80 17.34 39.73 38.71 67.63 38.69
Technological 37.77 27.63 16.34 41.68 35.15 68.29 37.81
Avg. 36.05 27.46 18.17 42.52 37.27 66.59 38.01
ibm-granite-granite-embedding-125m-english
Biological 65.94 61.42 50.92 62.06 46.86 71.23 59.74
Chemical 64.39 57.94 52.68 63.52 46.68 71.24 59.41
Environmental 65.87 60.32 47.60 64.84 52.70 73.30 60.77
Extraterrestrial 63.75 58.73 44.83 62.05 47.34 72.81 58.25
Geohazard 63.08 59.78 44.29 60.60 46.28 70.89 57.49
Meteorological&hydrological 65.83 61.04 45.60 62.71 45.98 71.02 58.70
Societal 66.36 65.53 44.21 62.98 50.31 71.93 60.22
Technological 63.95 63.52 47.03 63.49 48.21 73.06 59.88
Avg. 64.90 61.04 47.14 62.78 48.04 71.94 59.31
infly-inf-retriever-v1
Biological 76.15 68.85 71.20 68.79 52.90 78.79 69.45
Chemical 74.65 65.12 68.88 70.95 52.85 76.74 68.20
Environmental 73.45 66.91 69.84 70.05 56.45 77.55 69.04
Extraterrestrial 72.61 68.14 66.38 68.78 48.90 79.31 67.35
Geohazard 74.22 65.10 68.63 67.46 51.55 75.86 67.14
Meteorological&hydrological 74.21 69.31 70.86 68.37 52.17 76.76 68.61
Societal 75.49 70.65 69.64 68.07 57.23 78.49 69.93
Technological 73.03 68.48 69.23 68.84 52.73 78.12 68.41
Avg. 74.23 67.82 69.33 68.91 53.10 77.70 68.52
infly-inf-retriever-v1-1.5b
Biological 72.23 65.55 67.42 65.01 53.59 76.99 66.80
Chemical 71.18 63.09 65.04 67.03 51.93 75.97 65.71
Environmental 70.06 63.56 65.57 68.21 57.26 76.07 66.79
Extraterrestrial 67.37 61.83 63.32 65.93 51.12 77.02 64.43
Geohazard 71.49 62.52 65.70 65.56 52.29 75.73 65.55
Meteorological&hydrological 70.52 65.69 67.86 67.18 52.54 75.66 66.58
Societal 71.90 66.62 66.65 66.51 57.99 76.24 67.65
Technological 70.39 66.48 67.29 66.79 54.18 75.83 66.83
Avg. 70.64 64.42 66.11 66.53 53.86 76.19 66.29
Table 21: Performance of evaluated models under six search intents and eight event types under the exact search setting. Part III
QA QAdoc Twitter FC NLI STS Avg.
intfloat-e5-base-v2
Biological 67.85 62.93 58.90 62.08 47.03 74.82 62.27
Chemical 66.55 59.44 58.58 63.27 45.05 74.71 61.27
Environmental 67.45 63.11 60.69 64.54 49.66 74.99 63.41
Extraterrestrial 63.48 63.63 55.96 63.26 43.63 75.01 60.83
Geohazard 65.17 60.34 56.01 60.72 43.17 74.34 59.96
Meteorological&hydrological 66.14 62.64 59.13 62.21 44.88 74.08 61.52
Societal 64.92 67.05 59.00 60.68 46.63 74.02 62.05
Technological 64.58 65.38 58.65 61.44 44.27 75.15 61.58
Avg. 65.77 63.07 58.37 62.28 45.54 74.64 61.61
intfloat-e5-large-v2
Biological 59.31 63.11 56.68 62.40 51.02 75.08 61.27
Chemical 59.77 59.78 54.60 63.05 49.49 75.10 60.30
Environmental 62.61 63.25 57.49 64.79 54.51 75.21 62.98
Extraterrestrial 58.42 63.93 54.57 61.76 48.29 75.67 60.44
Geohazard 59.25 62.26 54.94 60.25 50.44 74.02 60.19
Meteorological&hydrological 60.07 63.55 57.59 62.98 49.77 75.12 61.51
Societal 60.04 66.56 56.92 60.85 53.09 74.75 62.04
Technological 61.74 64.95 56.80 62.20 51.07 74.97 61.95
Avg. 60.15 63.42 56.20 62.29 50.96 74.99 61.34
intfloat-e5-mistral-7b-instruct
Biological 18.64 19.90 22.40 32.79 22.57 43.06 26.56
Chemical 25.67 19.23 19.34 36.47 16.04 35.54 25.38
Environmental 24.96 17.65 21.97 28.46 22.05 41.30 26.06
Extraterrestrial 21.42 21.71 18.92 34.04 21.21 37.67 25.83
Geohazard 21.45 17.30 16.98 25.16 19.57 41.11 23.59
Meteorological&hydrological 20.67 15.84 20.54 24.92 20.85 37.62 23.41
Societal 21.53 24.83 16.94 31.62 23.46 35.64 25.67
Technological 18.86 19.62 18.73 34.85 17.34 44.61 25.67
Avg. 21.65 19.51 19.48 31.04 20.39 39.57 25.27
intfloat-e5-small-v2
Biological 66.38 63.19 61.91 61.38 48.09 75.04 62.67
Chemical 66.99 60.81 61.30 63.53 47.24 75.17 62.51
Environmental 67.61 64.30 62.38 65.73 50.80 75.57 64.40
Extraterrestrial 63.65 60.90 57.98 60.98 45.13 74.96 60.60
Geohazard 65.73 61.10 59.06 60.21 43.82 74.60 60.75
Meteorological&hydrological 66.89 63.88 61.80 62.53 46.50 74.76 62.73
Societal 64.53 65.51 61.41 60.10 47.63 74.69 62.31
Technological 65.19 64.31 61.32 61.88 47.51 73.84 62.34
Avg. 65.87 63.00 60.89 62.04 47.09 74.83 62.29
intfloat-multilingual-e5-large
Biological 68.63 62.63 64.36 60.27 49.91 74.90 63.45
Chemical 66.87 62.82 63.95 61.04 49.49 74.90 63.18
Environmental 66.74 63.98 64.06 62.20 54.65 74.76 64.40
Extraterrestrial 65.81 63.71 60.26 59.62 49.28 75.41 62.35
Geohazard 66.65 62.97 63.71 58.47 49.68 75.10 62.76
Meteorological&hydrological 68.57 63.87 63.41 60.45 50.40 74.53 63.54
Societal 67.61 67.61 64.44 58.97 54.07 74.86 64.59
Technological 67.47 66.39 65.80 61.08 50.71 76.02 64.58
Avg. 67.29 64.25 63.75 60.26 51.02 75.06 63.61
intfloat-multilingual-e5-large-instruct
Biological 70.03 63.24 64.67 66.51 50.24 64.07 63.13
Chemical 67.34 63.58 61.41 68.56 48.65 63.33 62.15
Environmental 68.93 64.34 63.38 69.14 50.64 62.96 63.23
Extraterrestrial 66.19 65.08 61.14 68.44 49.74 66.73 62.89
Geohazard 68.16 62.03 62.75 64.45 46.34 62.08 60.97
Meteorological&hydrological 68.59 66.05 63.42 66.52 50.51 64.22 63.22
Societal 69.62 68.27 63.92 66.58 51.19 65.31 64.15
Technological 68.25 66.61 65.26 67.51 47.72 65.81 63.53
Avg. 68.39 64.90 63.24 67.21 49.38 64.31 62.91
Table 22: Performance of evaluated models under six search intents and eight event types under the exact search setting. Part IV
QA QAdoc Twitter FC NLI STS Avg.
llmrails-ember-v1
Biological 64.90 62.60 49.37 60.23 45.66 74.47 59.54
Chemical 65.00 60.86 47.16 62.76 43.98 75.04 59.14
Environmental 65.68 61.87 47.82 62.25 48.17 74.18 59.99
Extraterrestrial 65.52 65.06 44.55 62.47 42.29 74.28 59.03
Geohazard 63.42 60.96 44.08 59.80 43.70 72.78 57.46
Meteorological&hydrological 64.02 64.12 45.55 60.74 42.42 72.11 58.16
Societal 65.11 66.53 43.32 59.41 46.03 74.16 59.09
Technological 64.87 64.73 45.61 60.78 44.23 74.90 59.19
Avg. 64.82 63.34 45.93 61.06 44.56 73.99 58.95
mixedbread-ai-mxbai-embed-large-v1
Biological 64.38 62.94 43.89 57.80 40.49 69.47 56.50
Chemical 64.29 60.87 42.16 59.77 39.28 70.09 56.08
Environmental 65.73 60.71 42.67 59.05 43.81 69.91 56.98
Extraterrestrial 65.50 65.11 38.38 60.94 38.15 69.16 56.21
Geohazard 62.01 60.48 40.11 56.35 40.05 67.06 54.34
Meteorological&hydrological 64.65 62.91 39.82 57.32 38.05 67.04 54.97
Societal 65.72 65.82 39.30 58.65 41.65 68.60 56.62
Technological 64.26 64.83 38.31 57.85 40.30 69.16 55.78
Avg. 64.57 62.96 40.58 58.47 40.22 68.81 55.93
nvidia-NV-Embed-v1
Biological 69.88 63.74 57.69 62.45 47.40 68.75 61.65
Chemical 69.31 56.86 57.02 59.12 45.99 66.66 59.16
Environmental 69.14 63.86 57.85 61.98 52.46 68.36 62.28
Extraterrestrial 65.64 62.43 53.94 57.93 45.98 69.16 59.18
Geohazard 66.26 62.29 55.90 58.56 48.33 65.33 59.44
Meteorological&hydrological 69.42 63.50 58.28 61.66 47.69 68.60 61.52
Societal 68.68 66.72 57.45 58.88 50.42 67.42 61.59
Technological 68.59 64.78 56.57 59.77 48.24 68.62 61.10
Avg. 68.36 63.02 56.84 60.04 48.31 67.86 60.74
nvidia-NV-Embed-v2
Biological 76.42 69.25 42.42 69.75 58.26 77.86 65.66
Chemical 74.82 69.52 42.07 68.68 57.96 76.22 64.88
Environmental 75.31 68.66 43.82 69.63 61.09 77.21 65.95
Extraterrestrial 73.06 70.08 44.88 68.30 56.31 78.18 65.14
Geohazard 73.91 67.35 41.35 66.72 57.88 75.91 63.85
Meteorological&hydrological 75.33 69.91 43.11 68.37 58.88 76.85 65.41
Societal 75.38 72.11 44.57 68.78 61.06 76.54 66.41
Technological 73.91 71.01 43.18 68.85 58.42 77.31 65.45
Avg. 74.77 69.74 43.18 68.64 58.73 77.01 65.34
sentence-transformers-all-mpnet-base-v2
Biological 14.00 7.64 17.18 43.65 27.58 35.38 24.24
Chemical 13.09 9.93 14.53 48.81 24.14 35.84 24.39
Environmental 12.45 10.01 18.17 47.32 30.32 35.48 25.62
Extraterrestrial 16.24 8.11 14.82 48.79 27.48 42.02 26.24
Geohazard 14.96 7.75 13.30 43.82 27.70 36.33 23.98
Meteorological&hydrological 16.52 7.84 16.39 44.34 28.07 36.99 25.02
Societal 20.10 14.49 18.41 45.98 27.31 39.06 27.56
Technological 13.16 12.41 16.55 47.68 26.44 36.75 25.50
Avg. 15.06 9.77 16.17 46.30 27.38 37.23 25.32
thenlper-gte-base
Biological 10.09 5.90 45.04 61.49 44.70 49.42 36.11
Chemical 9.73 5.62 40.67 62.27 41.06 53.75 35.52
Environmental 6.03 4.75 38.62 63.00 48.95 40.21 33.59
Extraterrestrial 10.42 4.54 33.58 59.45 41.39 44.64 32.34
Geohazard 7.96 3.33 34.20 58.54 41.68 43.00 31.45
Meteorological&hydrological 9.16 5.31 38.46 60.76 40.55 43.60 32.97
Societal 10.00 7.88 39.76 60.39 44.40 48.17 35.10
Technological 10.39 5.44 38.00 60.50 40.08 50.33 34.13
Avg. 9.22 5.35 38.54 60.80 42.85 46.64 33.90
Table 23: Performance of evaluated models under six search intents and eight event types under the exact search setting. Part V

### v2

@@
 Disaster Management
-Kai Yin1  Xiangjue Dong1  Chengkai Liu1  Lipai Huang1 Yiming Xiao1Zhewei Liu2Ali Mostafavi1James Caverlee1 1Texas A&M University  2University of Toronto {kai_yin,xj.dong,liuchengkai,lipai.huang,yxiao,mostafavi,caverlee}@tamu.edu zwei.liu@utoronto.ca
+Kai Yin1   Xiangjue Dong1,*   Chengkai Liu1,*   Lipai Huang1 Yiming Xiao1Zhewei Liu2Ali Mostafavi1James Caverlee1 1Texas A&M University  2University of Toronto {kai_yin,lipai.huang,yxiao,mostafavi,caverlee}@tamu.edu zwei.liu@utoronto.ca *Corresponding Authors: xj.dong@tamu.edu, liuchengkai@tamu.edu
 1 Introduction
@@
 $~~~~~~~~~~~\quad\longrightarrow\;(q,\;psg)~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~(1)$
-where $LLM_{info}$ and $LLM_{query}$ are LLMs prompted to generate retrieval information needs statements and the query-passage pair respectively, $P_{IN}$ and $P_{QG}$ are prompts for information needs and query generation, $P_{seed}$ is the domain passage, $q$ is the synthesized user query, and $psg$ is the corresponding relevant passage.
+where $LLM_{info}$ and $LLM_{query}$ are LLMs prompted to generate retrieval information needs statements and the query-passage pair respectively, $P_{IN}$ and $P_{QG}$ are prompts for information needs and query generation, $P_{seed}$ is the domain passage, $q$ is the synthesized user query, and psgpsg is the corresponding relevant passage.
 To ensure generated queries align with the core characteristics and objectives of each search intent, we design intent-specific prompts for both stages of query generation.
-The full prompt templates for each intent are provided in Appendix B.
+The full prompt templates for each intent are provided in Appendix C.
 For each search task, we generate 200 unique user queries by prompting an LLM with randomly sampled domain-specific passages, resulting in 9,600 queries.
@@
 For each search intent, relevance is defined to align with its specific objectives, reflecting the varying interpretations of “relevance” across different task types (Dai et al., 2022).
-Full prompt templates are provided in Appendix C.
+Full prompt templates are provided in Appendix D.
 Relevance scores are assigned on a 4-point scale (0 to 3) for all intents, except STS, which follows a 6-level scale as in Agirre et al.
@@
 Name Size Bin QA QAdoc TW FC NLI STS Avg Avg (%)
-inf-retriever-v1 7B XL 74.23 67.82 69.33 68.91 53.10 77.70 68.52 66.90 2.36
-SFR-Embedding-Mistral 7B XL 71.76 67.58 70.42 70.62 50.86 73.61 67.47 66.75 1.08
-inf-retriever-v1-1.5b 1.5B XL 70.64 64.42 66.11 66.53 53.86 76.19 66.29 65.54 1.13
-NV-Embed-v2 7B XL 74.77 69.74 43.18 68.64 58.73 77.01 65.34 64.45 1.36
-multilingual-e5-large 560M Large 67.29 64.25 63.75 60.26 51.02 75.06 63.61 62.08 2.41
-multilingual-e5-large-instruct 560M Large 68.39 64.90 63.24 67.21 49.38 64.31 62.91 62.06 1.34
-e5-small-v2 33M Small 65.87 63.00 60.89 62.04 47.09 74.83 62.29 60.91 2.22
-e5-base-v2 109M Medium 65.77 63.07 58.37 62.28 45.54 74.64 61.61 60.23 2.25
-e5-large-v2 335M Large 60.15 63.42 56.20 62.29 50.96 74.99 61.34 60.20 1.85
-NV-Embed-v1 7B XL 68.36 63.02 56.84 60.04 48.31 67.86 60.74 59.24 2.47
-granite-embedding-125m 125M Medium 64.90 61.04 47.14 62.78 48.04 71.94 59.31 58.67 1.06
-llmrails-ember-v1 335M Large 64.82 63.34 45.93 61.06 44.56 73.99 58.95 58.35 1.02
-arctic-embed-m-v2.0 305M Medium 61.44 62.50 47.80 58.04 42.34 65.28 56.23 54.75 2.64
-mxbai-embed-large-v1 335M Large 64.57 62.96 40.58 58.47 40.22 68.81 55.93 55.42 0.92
-gte-base-en-v1.5 137M Medium 60.66 56.06 47.06 52.49 39.88 71.27 54.57 53.71 1.57
-gte-large-en-v1.5 434M Large 67.70 58.55 40.31 53.09 34.81 67.30 53.63 53.09 1.01
-snowflake-arctic-embed-l-v2.0 568M Large 55.34 59.43 38.75 60.42 41.19 63.38 53.09 51.90 2.23
-bge-base-en-v1.5 109M Medium 49.67 53.97 47.03 58.47 35.58 64.43 51.52 50.71 1.57
-bge-large-en-v1.5 335M Large 57.02 54.72 32.71 55.22 35.19 65.14 50.00 48.99 2.03
-gte-Qwen2-7B-instruct 7B XL 62.47 39.12 41.40 29.98 47.28 70.98 48.54 47.07 3.03
-bge-small-en-v1.5 33M Small 47.38 45.51 27.82 52.13 27.20 56.86 42.82 42.07 1.75
-snowflake-arctic-embed-m-v1.5 109M Medium 25.73 30.56 18.31 48.24 43.04 64.95 38.47 36.25 5.79
-snowflake-arctic-embed-s 33M Small 36.05 27.46 18.17 42.52 37.27 66.59 38.01 33.59 11.63
-snowflake-arctic-embed-l 335M Large 40.82 30.41 15.32 32.70 34.62 56.82 35.11 31.35 10.72
-Linq-Embed-Mistral 7B XL 35.61 31.41 26.55 39.46 27.30 45.69 34.34 33.52 2.38
-thenlper-gte-base 109M Medium 9.22 5.35 38.54 60.80 42.85 46.64 33.90 31.78 6.24
-snowflake-arctic-embed-m 109M Medium 33.35 14.25 8.64 35.30 38.93 56.88 31.23 28.83 7.67
-all-mpnet-base-v2 109M Medium 15.06 9.77 16.17 46.30 27.38 37.23 25.32 25.05 1.08
-e5-mistral-7b-instruct 7B XL 21.65 19.51 19.48 31.04 20.39 39.57 25.27 24.05 4.82
-gte-Qwen2-1.5B-instruct 1.5B XL 13.98 22.21 19.61 23.90 18.00 31.20 21.48 21.27 0.98
+Linq-Embed-Mistral 7B XL 74.40 70.50 64.22 70.77 52.56 71.35 67.30 66.98 0.48
+SFR-Embedding-Mistral 7B XL 71.50 67.34 69.62 70.39 51.08 72.71 66.71 66.39 0.48
+inf-retriever-v1 7B XL 72.84 66.92 66.37 65.76 52.02 76.00 66.65 65.98 1.01
+inf-retriever-v1-1.5b 1.5B XL 69.47 64.40 63.08 65.49 54.14 73.96 65.09 64.85 0.37
+NV-Embed-v2 7B XL 74.55 69.51 42.55 68.39 58.39 76.13 64.92 64.57 0.54
+gte-Qwen2-1.5B-instruct 1.5B XL 69.96 59.21 65.21 62.84 55.73 73.61 64.43 64.24 0.29
+multilingual-e5-large 560M Large 67.08 64.08 62.99 60.06 51.20 74.14 63.26 62.79 0.74
+e5-mistral-7b-instruct 7B XL 65.65 65.16 63.42 67.94 47.68 66.39 62.71 61.99 1.15
+multilingual-e5-large-instruct 560M Large 68.14 64.72 62.46 66.96 48.75 63.53 62.43 62.01 0.67
+e5-small-v2 33M Small 65.66 62.84 60.10 61.78 47.12 73.93 61.90 61.48 0.68
+e5-base-v2 109M Medium 65.54 62.91 57.76 62.11 45.52 73.73 61.26 60.72 0.88
+e5-large-v2 335M Large 60.03 63.24 55.48 62.03 50.96 74.09 60.97 60.45 0.85
+NV-Embed-v1 7B XL 68.14 62.87 56.13 59.85 48.25 67.11 60.39 59.60 1.31
+granite-embedding-125m 125M Medium 64.63 60.85 46.55 62.56 48.11 71.06 58.96 58.60 0.61
+gte-Qwen2-7B-instruct 7B XL 70.30 47.65 63.24 31.87 53.88 74.86 56.97 55.99 1.72
+snowflake-arctic-embed-m-v2.0 305M Medium 61.28 62.31 47.20 57.84 42.43 64.56 55.94 55.15 1.41
+mxbai-embed-large-v1 335M Large 64.37 62.79 40.07 58.30 40.26 67.96 55.62 55.25 0.67
+gte-base-en-v1.5 137M Medium 60.46 55.85 46.44 52.34 39.85 70.41 54.22 53.93 0.53
+bge-base-en-v1.5 109M Medium 51.65 52.89 46.78 60.13 41.41 68.56 53.57 53.13 0.82
+gte-large-en-v1.5 434M Large 67.46 58.37 39.71 52.90 34.79 66.51 53.29 53.21 0.15
+snowflake-arctic-embed-l-v2.0 568M Large 55.20 59.29 38.26 60.23 41.23 62.64 52.81 52.32 0.93
+bge-large-en-v1.5 335M Large 56.88 54.56 32.32 55.03 35.25 64.43 49.74 49.04 1.41
+bge-small-en-v1.5 33M Small 56.87 51.24 25.19 55.30 32.95 64.46 47.67 47.00 1.41
+snowflake-arctic-embed-s 33M Small 38.69 28.82 21.43 47.30 40.02 66.95 40.54 38.15 5.90
+snowflake-arctic-embed-m-v1.5 109M Medium 25.66 30.43 18.09 48.10 42.98 64.20 38.24 36.85 3.63
+snowflake-arctic-embed-l 335M Large 40.73 30.33 15.11 32.60 34.44 56.11 34.89 32.17 7.80
+thenlper-gte-base 109M Medium 9.16 5.34 38.06 60.58 42.80 45.99 33.66 32.22 4.28
+snowflake-arctic-embed-m 109M Medium 33.26 14.22 8.62 35.16 38.75 56.21 31.02 29.42 5.16
+snowflake-arctic-embed-m-long 137M Medium 21.43 10.84 19.49 36.20 41.90 55.00 30.81 29.30 4.90
+thenlper-gte-small 33M Small 18.20 9.08 11.04 49.81 37.71 55.47 30.22 29.43 2.61
 Table 3: Performances of 30 evaluated IR models in DisastIR.
 Models are ranked by their overall performance under exact search (highest to lowest) in DisastIR.
-“Size Bin” indicates its model parameter size bin category (small, medium, large, and extra large as defined in Appendix G).
+“Size Bin” indicates its model parameter size bin category (small, medium, large, and extra large as defined in Appendix H).
 “TW” represents Twitter.
@@
 Bold indicates the highest value, and underline indicates the second-highest.
+E5-small, base, large-v2, and granite-embedding use knowledge distillation during fine-tuning, which involves additional training signals.
+Performances across different event types are shown in Table 4.
 4 DisastIR Benchmark Analysis
@@
 This wide distribution captures the diversity of disaster management-related texts, including both brief updates and detailed descriptions like event summaries or emergency protocols.
+The corpus comprises both original and synthetic passages, with synthetic passages making up only 6.8 % of the total.
+They come from two sources: 8,000 passages generated from page content ( 3.3 %) and 8,464 document-level summaries serving as proxies for full documents ( 3.5 %).
+Synthetic passages are introduced to enhance diversity and support varied search intents.
+Original passages, extracted from PDFs, are formal in style and, when chunked, often miss the broader document context.
+Some search scenarios, such as Twitter retrieval, require informal text featuring emojis, hashtags, or colloquial expressions, which are largely absent in PDF text.
+Other scenarios, such as QAdoc, demand whole-document understanding, where LLM-generated summaries provide an effective substitute since full texts typically exceed the input limits of IR models.
+Evaluation of LLM-generated passages is validated in Appendix B.2.
 Labeled Query-Passage Pairs.
@@
 This larger pool in Twitter tasks suggests greater divergence in model outputs, indicating lower agreement among retrieval models when ranking passages in social media contexts within disaster management scenarios.
-Additional analyses of labeled query-passage pairs are provided in Appendix D.
+Additional analyses of labeled query-passage pairs are provided in Appendix E.
 4.2 LLM-based vs. Human Labeling
@@
 MS MARCO (Bajaj et al., 2016) and TriviaQA (Joshi et al., 2017) are for QA, ALLNLI (sentence-transformers, 2021) and XNLI (Conneau et al., 2018) for NLI, Climate-Fever (Diggelmann et al., 2021) for FC, and STSB (Cer et al., 2017) for STS.
-Appendix E provides details on the construction of LVHL.
+Appendix F provides details on the construction of LVHL.
 The LLM-based relevance scores for each query-passage pair in LVHL are computed as described in Section 3.6.
@@
 These suggest that LLM-generated relevance scores align well with human judgments and can reliably substitute for manual annotation in DisastIR.
+We further conduct a controlled in-domain annotation experiment.
+Specifically, we sample 96 query–passage pairs from the DisastIR labeling pool, covering all 48 intent–event type combinations, and ask three PhD students specializing in disaster management to independently annotate them using the same multi-level relevance scale as in our LLM pipeline.
+Annotation guidelines are adapted from the chain-of-thought reasoning procedure employed for LLM labeling.
+Inter-annotator agreement was substantial, with Fleiss’ kappa reaching 0.777.
+Comparing LLM-generated scores with human annotations yielded Cohen’s kappa values of 0.681–0.734 across individual annotators, 0.690 when averaged, and 0.803/0.715 under majority vote (minimum/maximum tie-breaking), again demonstrating substantial agreement (Table 6).
+These findings confirm that LLM-based labeling is not only consistent with external human judgments (as shown in LVHL) but also robustly aligned with expert assessments within the disaster management domain.
 4.3 LLM vs. Human-generated User Query
@@
 All query-passage pairs are annotated using the same method as in DisastIR.
-Appendix F provides full details on the construction of LVHQ.
+Appendix G provides full details on the construction of LVHQ.
 We evaluate all selected baseline models using LVHQ under exact search for both human- and LLM-generated queries (see Section 5 for evaluation setup up).
 Model performance, measured by NDCG@10, shows highly consistent results across the two query types, with a Kendall’s $\tau$ of 0.9264, indicating strong agreement in model evaluations.
+Figure 3: Distribution of evaluated models’ performances across all 48 tasks. The full name of each model in the X axis is listed in Model Name column in Table 3.
 5 Experimental Setup
@@
 DisastIR is adopted to comprehensively evaluate open-source IR models and support the selection of suitable IR models for real-world disaster management applications.
-Models are chosen based on three criteria: (1) strong performance on the MTEB retrieval benchmark; (2) inclusion in widely adopted embedding families such as BGE (Chen et al., 2024; Xiao et al., 2024), E5 (Wang et al., 2022, 2023), Snowflake Arctic (Merrick, 2024), and GTE (Li et al., 2023; Zhang et al., 2024a), which are commonly used as baselines and in downstream IR tasks (Sun et al., 2024; Xu et al., 2024; Lee et al., 2024b, a; Cao, 2025; Park et al., 2025); and (3) coverage of different types of search tasks adapted from Wang et al.
-(2024b).
-We select 30 models with parameter sizes ranging from 33 million to 7 billion. Detailed descriptions of these models and their implementations are provided in Appendix G.
-Figure 3: Distribution of evaluated models’ performances across all 48 tasks. The full name of each model in the X axis is listed in Model Name column in Table 3.
+Models are chosen based on two criteria: (1) strong performance on the MTEB retrieval benchmark; and (2) inclusion in widely adopted embedding model families such as BGE (Chen et al., 2024; Xiao et al., 2024), E5 (Wang et al., 2022, 2023), Snowflake Arctic (Merrick, 2024), and GTE (Li et al., 2023; Zhang et al., 2024a), which are commonly used as baselines and in downstream IR tasks (Sun et al., 2024; Xu et al., 2024; Lee et al., 2024b, a; Cao, 2025; Park et al., 2025).
+We select 30 models with parameter sizes ranging from 33 million to 7 billion. Detailed descriptions of these models and their implementations are provided in Appendix H.
 5.2 Evaluation
-We evaluate model performance under two retrieval settings – exact and ANN – using Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) as the primary metric, consistent with prior works.
+We evaluate model performance under two retrieval settings, exact and ANN, using Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) as the primary metric, consistent with prior works.
 (1) Exact Brute-force Retrieval.
@@
 To reflect real-world large-scale disaster information retrieval scenarios, we also evaluate model performance during the candidate generation stage using ANN search.
-We adopt Usearch (Vardanian, 2023), a high-performance, memory-efficient library based on the HNSW (hierarchical navigable small world) algorithm (Malkov and Yashunin, 2018), to retrieve top $k$ passages per query using precomputed embeddings.
+We adopt the HNSW (hierarchical navigable small world) algorithm (Malkov and Yashunin, 2018), to retrieve top $k$ passages per query using precomputed embeddings.
 For fair comparison, $k$ is set to match the value used in exact search.
+Figure 4: Best-performing models in each search task.
 Figure 5: Comparison between DisastIR and MTEB model rankings. Legend shapes indicate the model size bin: $\blacklozenge$ XL, $\blacktriangle$ Large, $\blacksquare$ Medium, $\bullet$ Small.
@@
 6.1 Overall Performance
-Table 3 summarizes the overall performance of all 30 evaluated models across all queries in DisastIR, with detailed results for each search task provided in Appendix H.
-The inf-retriever-v1 model ranks highest in both exact and ANN search settings, followed closely by SFR-Embedding-Mistral (1.52% and 0.22% lower, respectively).
-Among non-XL models, multilingual-e5-large performs best, reaching 92.83% and 92.79% of the top model’s performance.
-Notably, the lightweight e5-small-v2 model (33M parameters) achieves 90.91% and 91.05% of the top model’s performance, despite being 212 times smaller in size.
-The Snowflake-arctic-embed-s model shows the largest performance drop (11.63%) under ANN search compared to exact search (Table 3).
-Most models exhibit drops within 5%; only five exceeded this margin, four of which belong to snowflake family, indicating strong robustness when switching from exact to ANN search in DisastIR.
-Among the top-performing models analyzed above, SFR-Embedding-Mistral has the smallest drop (1.08%), while multilingual-e5-large has the largest (2.41%).
+Table 3 summarizes the overall performance of all 30 evaluated models across all queries in DisastIR, with detailed results for each search task provided in Appendix I.
+The Linq-Embed-Mistral model achieves the best performance under exact and ANN search settings, followed closely by SFR-Embedding-Mistral (0.877% and 0.881% lower, respectively).
+Among all non-XL models, multilingual-e5-large performs best, reaching 94.0% and 93.00% of the top model’s performance.
+Notably, the lightweight e5-small-v2 model (33M parameters) achieves 91.98% and 91.17% of the top model’s performance, despite being 212 times smaller in size.
+The E5-V2 series (Wang et al., 2022) and Granite-Embedding (Awasthy et al., 2025) leverage knowledge distillation during fine-tuning, achieving performance that surpasses all models of the same scale and even outperforms many substantially larger models.
+This highlights the effectiveness of knowledge distillation in enhancing the performance of smaller models.
+The Snowflake-arctic-embed-l model shows the largest performance drop (7.80%) under ANN search compared to exact search (Table 3).
+Most models exhibit drops within 2%; only five exceeded this margin, four of which belong to snowflake family, indicating strong robustness when switching from exact to ANN search in DisastIR.
 All subsequent analyses are based on exact search; analyses under ANN search can be conducted similarly.
-Figure 5: Best-performing models in each search task.
 6.2 Performance across all 48 Tasks
 Figure 3 presents the performance distribution of all evaluated models across all 48 search tasks.
-While inf-retriever-v1 achieves the highest average NDCG@10, its median performance is lower than that of SFR-Embedding-Mistral, and it exhibits greater variability across tasks, as reflected by a larger interquartile range (IQR).
-This suggests that inf-retriever-v1 is less stable across diverse search tasks in DisastIR.
+All top-5 performance models show great variability across tasks, as reflected by the large interquartile range (IQR).
+This highlights the limited cross-task robustness of current general domain retrieval models and underscores the need to design methods that enhance cross-task consistency, rather than optimizing solely for higher average performance.
 As shown in Figure 4, no single model consistently outperforms others across all 48 tasks.
-Instead, top performance is distributed among three models: inf-retriever-v1, SFR-Embedding-Mistral, and NV-Embed-v2.
-inf-retriever-v1 achieves the best results in only nine tasks, primarily in STS.
+Instead, top performance is distributed among four models: Linq-Embed-Mistral, inf-retriever-v1, SFR-Embedding-Mistral, and NV-Embed-v2.
 This highlights the complexity and diversity of disaster management-related retrieval tasks and reinforces the need for domain-specific IR models in real-world disaster management scenarios.
-Appendix I provides additional analyses of model performance across 48 tasks.
+Appendix J provides additional analyses of model performance across 48 tasks.
+Among these top-performing models, only NV-Embed-v2 is accompanied by a public technical report (Lee et al., 2024a), enabling a closer look at its varying performance across intents.
+Its training data emphasizes fact-checking, NLI, and QA but excludes Twitter, explaining strong results on QA and NLI and weaker performance on Twitter.
+Template usage further matters: NV-Embed-v2 benefits from alignment between training and inference for QA, QAdoc, FC, and NLI, but lacks such alignment for Twitter.
+Architecturally, unlike models relying on the final <EOS> token, NV-Embed-v2 employs a latent attention layer, which helps it achieve the highest average NDCG@10 (69.39) among top models when Twitter is excluded.
 6.3 Comparison with General Domain
@@
 Ranking value of each model is based on overall performance in DisastIR and official retrieval scores from the MTEB English leaderboard.
-The Spearman correlation between the two rankings is 0.225 ($p=0.251$), indicating no significant correlation.
+The Spearman correlation between the two rankings is 0.252 ($p=0.188$), indicating no significant correlation.
 This suggests that strong performance on general-domain benchmarks does not guarantee effectiveness in disaster management-related retrieval.
-For example, models like Linq-Embed-Mistral and snowflake-arctic-embed-l perform well in MTEB but poorly in DisastIR, while models from the E5 family show the opposite trend.
-Although inf-retriever-v1 ranks highest in both DisastIR and MTEB, detailed analysis in Section 6.2 show that it achieves top performance in fewer than 20% of DisastIR tasks, mostly those related to STS.
+For example, models in snowflake family perform well in MTEB but poorly in DisastIR, while models from the E5 family show the opposite trend.
 Furthermore, when computational resources are limited and large models are impractical to serve, relying solely on MTEB rankings for model selection, such as choosing snowflake-arctic-embed-l, may fail to retrieve critical or relevant content.
@@
 Acknowledgments
-Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing
+This work used ACES at TAMU, DeltaAI and Delta GPU at the National Center for Supercomputing Applications through allocation CIV250019 and CIV250021 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S.
+National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
 Appendix A Structural PDF File Processing Pipeline
+Model Bio Chem Env Extra Geo MH Soc Tech
+Linq-Embed-Mistral 68.07 66.75 67.82 67.30 65.40 67.13 68.56 67.36
+SFR-Embedding-Mistral 67.74 66.70 68.03 66.43 65.40 67.03 68.15 67.37
+inf-retriever-v1 67.75 66.40 66.92 65.04 65.52 66.66 68.15 66.75
+inf-retriever-v1-1.5b 65.71 64.45 65.36 63.33 64.31 65.45 66.59 65.54
+NV-Embed-v2 65.23 64.51 65.55 64.79 63.51 64.94 65.95 64.89
+gte-Qwen2-1.5B-instruct 65.15 64.35 64.94 62.56 63.16 63.97 66.12 65.16
+multilingual-e5-large 63.00 62.91 64.01 62.11 62.59 63.10 64.26 64.10
+e5-mistral-7b-instruct 63.33 62.13 63.49 62.25 60.88 62.70 63.85 63.04
+multilingual-e5-large-instruct 62.64 61.79 62.76 62.48 60.49 62.70 63.60 62.95
+e5-small-v2 62.73 62.56 64.17 60.59 60.71 62.68 62.29 62.29
+e5-base-v2 61.86 61.34 63.05 60.97 61.08 61.96 62.98 61.75
+e5-large-v2 60.09 59.49 60.84 59.77 59.18 60.48 61.15 60.74
+NV-Embed-v1 61.37 58.86 61.91 58.84 59.13 61.13 61.23 60.66
+granite-embedding-125m 58.00 57.52 58.09 57.21 56.80 57.89 58.32 57.61
+arctic-embed-m-v2.0 56.47 55.78 56.14 55.05 54.88 55.66 55.98 55.29
+mxbai-embed-large-v1 56.18 55.38 55.96 54.77 54.56 55.47 55.88 55.07
+gte-base-en-v1.5 54.78 54.29 55.49 53.19 52.99 54.19 54.60 53.98
+bge-base-en-v1.5 53.57 52.78 54.18 52.38 51.98 53.18 53.77 52.88
+gte-large-en-v1.5 63.23 62.54 63.86 62.32 61.92 63.05 64.10 63.07
+snowflake-arctic-embed-l-v2.0 53.86 53.09 54.44 52.78 52.36 53.58 53.96 53.17
+bge-large-en-v1.5 53.18 52.49 53.88 52.08 51.78 52.79 53.49 52.58
+gte-Qwen2-7B-instruct 52.90 52.20 53.70 51.90 51.60 52.60 53.20 52.30
+bge-small-en-v1.5 45.05 44.45 35.75 44.91 44.14 45.23 45.95 44.73
+snowflake-arctic-embed-s 40.91 40.51 43.60 38.26 38.71 40.15 41.65 40.50
+snowflake-arctic-embed-m-v1.5 42.30 42.79 44.14 40.49 41.20 41.99 42.47 42.19
+snowflake-arctic-embed-l 35.15 35.05 37.77 34.28 32.53 33.81 35.92 34.59
+thenlper-gte-base 35.84 35.28 33.35 32.19 31.19 32.71 34.83 33.84
+snowflake-arctic-embed-m 30.72 31.00 32.51 31.40 29.45 30.71 31.62 30.74
+snowflake-arctic-embed-m-long 30.82 30.67 32.16 30.77 29.44 30.00 31.87 30.75
+thenlper-gte-small 31.92 31.53 28.83 30.25 27.72 30.26 29.96 31.27
+Table 4: Performance of 30 open-source retrieval models under the exact-search setting across different disaster event types.
+Each cell shows the mean NDCG@10 over the 6 search tasks for that event type.
 Figure 6: Cohen’s kappa scores between LLM-based and human-annotated relevance labels across all LVHL datasets, as described in Section 4.2
@@
 Hence, explicit consent was not required.
+We chose PDFs because they typically contain more structured, information-rich, and credible content, often originating from peer-reviewed publications or official institutions.
 PDF files are collected using googlesearch-python (v1.3.0) and processed with PyMuPDF (v1.24.10) for content extraction.
@@
 An ANN index is built to retrieve the top-$k$ nearest chunks, and pairs with cosine similarity above 0.9 are removed.
-Appendix B Prompt Templates for Query Generation
-Prompts for query generation based on disaster management-related passage under different search intents for QA, QAdoc, Twitter, FC, NLI, STS are in Tables 4, 5, 6, 7, 8, and 9.
-Appendix C Prompt Templates for Relevance Labeling
+Since our benchmark emphasizes retrieval performance across event types and search intents rather than source analysis, we summarize the corpus by the number of PDFs per general event type.
+Each general event type includes multiple specific event types, which were used as search queries.
+The number of collected PDFs is thus correlated with the number of associated specific types, as shown in Table 5.
+General event type Number of PDFs (percentage)
+MH 1683 (19.89%)
+Bio 2343 (27.68%)
+Geo 868 (10.26%)
+Tech 1598 (18.88%)
+Extra 254 (3.00%)
+Chem 767 (9.06%)
+Soc 234 (2.76%)
+Env 717 (8.47%)
+Table 5: Distribution of event types by number of PDFs
+Appendix B Evaluation of LLM-generated passages
+B.1 Evaluation of LLM document summaries
+Experiment design.
+We evaluate whether each LLM-generated summary accurately captured the main content of the original material, which motivated their inclusion in the corpus.
+A total of 48 summaries (six per general event type) are randomly sampled.
+Three PhD students independently annotate each summary for fluency and content accuracy, following detailed guidelines adapted from Song et al.
+(2024).
+Results and analysis.
+Fluency is defined as the percentage of summaries judged grammatically and semantically well-formed, while content accuracy measures whether summaries captured the main ideas of the source material.
+The average results show 100% fluency and 96.88% content accuracy.
+The Fleiss’ Kappa score of 0.793 indicated substantial inter-annotator agreement, supporting the reliability of the evaluation.
+Overall, LLM-generated summaries are fluent and highly faithful to the original documents.
+B.2 Evaluation of LLM-generated passages from page content
+Experiment design.
+We further assess LLM-generated passages created directly from page content, focusing on their fluency and adherence to style requirements.
+The QAdoc intent is excluded, as its format (document summaries) had already been validated.
+We randomly sampled 40 passages (eight per intent across five intents) and asked three PhD students to evaluate them.
+The evaluators followed the same style guidelines provided to the LLM during generation.
+Results and analysis.
+LLM-generated passages achieved an average fluency score of 98.75% and a style compliance score of 92.5%.
+The Fleiss’ Kappa of 0.760 again confirmed substantial agreement among annotators.
+These results indicate that the generated passages are both fluent and well aligned with intent-specific stylistic requirements.
+B.3 Role of noisy passages in realistic IR evaluation
+While our evaluations confirm the overall quality of LLM-generated content, we emphasize the value of including imperfect or noisy passages in the corpus.
+Real-world IR corpora naturally contain irrelevant or off-topic data, and such noise can enhance the realism of evaluations.
+Introducing noisy passages allows us to test the robustness of IR models by assessing whether they can correctly identify irrelevant content and assign low relevance scores.
+For example, a generated tweet unrelated to a query serves as a negative case; an effective IR model should detect this mismatch and rank it accordingly.
+Appendix C Prompt Templates for Query Generation
+Prompts for query generation based on disaster management-related passage under different search intents for QA, QAdoc, Twitter, FC, NLI, STS are in Tables 7, 8, 9, 10, 11, and 12.
+Appendix D Prompt Templates for Relevance Labeling
 This section presents the prompt templates used for LLM-based relevance judgments across six search intents, employing three prompting strategies: Zero-shot Direct Scoring, Chain-of-Thought Decomposed Reasoning, and Multi-Dimensional Attribute Scoring.
@@
 (2024); Farzi and Dietz (2024), as shown in Table 13.
-Based on these templates, we design relevance prompts for FC, NLI, and STS tasks, shown in Tables 11, 12, and 13, respectively.
+Based on these templates, we design relevance prompts for FC, NLI, and STS tasks, shown in Tables 14, 15, and 16, respectively.
 For STS, we adopt only Zero-shot Direct Scoring, as our preliminary experiments show it yields higher agreement with human labels (Cohen’s kappa).
 The estimated cost of generating 9,600 user queries and labeling over 1.3 million query-passage using GPT-4o-mini API is about \$1,400.
-Appendix D Additional Analyses of Labeled query-passage pairs
+Appendix E Additional Analyses of Labeled query-passage pairs
 As shown in Table 17, in certain retrieval tasks, such as NLI_Bio, NLI_Geo, the number of query-passage pairs assigned the highest relevance score is smaller than the number of user queries.
@@
 (2024b).
-Appendix E LVHL Dataset Construction
+Appendix F LVHL Dataset Construction
 We use the names of 301 specific disaster event types as queries to search for disaster management-related user queries within each selected open-source dataset listed in Table 19.
@@
 It is not suitable for benchmarking retrieval models in the disaster management area, as most queries are drawn from training sets of the source datasets.
-Appendix F LVHQ Dataset Construction
+Appendix G LVHQ Dataset Construction
 We sample 48 domain passages developed in Section 3.3, ensuring one passage per retrieval task and keeping all sampled passages different from those used in developing DisastIR.
 For each passage, a domain expert in the disaster management field is asked to read the passage and write a realistic user query that reflects a practical information need based on the content, resulting in 48 human-authored queries.
-The Human expert is given the same instructions for the query written (shown in Tables 4, 5, 6, 7, 8, 9) as those given to LLM to ensure fair comparison.
+The Human expert is given the same instructions for the query written (shown in Tables 7, 8, 9, 10, 11, 12) as those given to LLM to ensure fair comparison.
 In parallel, for the same set of passages, we also generate 48 queries using LLM in the same way as described in Section 3.4.
@@
 Figure 8: Distribution of outliers of evaluated models’ performances
-Appendix G Information of Evaluated Models and Model Implementation
+Appendix H Information of Evaluated Models and Model Implementation
 Detailed information on all selected models is summarized in Table 20.
@@
 All models are run on a single NVIDIA A6000 GPU using HuggingFace Transformers, following the configurations specified in the official implementations.
-Appendix H Performance of Evaluated Models
-Performance of all evaluated models in all 48 search tasks in DisastIR is shown in Tables 19, 20, 21, 22, and 23.
-Appendix I Additional Analyses of Model Performance across 48 Tasks
-NV-Embed-v2 achieves the best performance on all QAdoc and NLI-related tasks (See Table 3 and Figure 5 in the main content).
+Appendix I Performance of Evaluated Models
+Performance of all evaluated models in all 48 search tasks in DisastIR is shown in Tables 22, 23, 24, 25, and 26.
+Appendix J Additional Analyses of Model Performance across 48 Tasks
+NV-Embed-v2 achieves the best performance on all NLI-related tasks (See Table 3 and Figure 4 in the main content).
 However, as shown in Figure 7, its poor results on Twitter-related tasks significantly lower its overall performance in DisastIR.
-This reflects its limitation in handling the informal, noisy, and contextually ambiguous nature of social media content.
+This reflects its limitation in handling informal, noisy, and contextually ambiguous nature of social media content.
 Given the importance of Twitter as a real-time, crowd-sourced information source during disasters (Alam et al., 2021; Yin et al., 2024; Lei et al., 2025), this weakness raises concerns about its reliability in real-world disaster response scenarios.
-All three models perform poorly on NLI-related tasks, with the best achieving only an average score of 58.73 (Figure 7).
+All four models perform poorly on NLI-related tasks, with the best achieving only an average score of 58.39 (Figure 7).
 Further analysis of outliers in the box plot (See Figure 3 in the main content) reveals that tasks causing significant performance drops consistently involve NLI search intents (Figure 8).
@@
 Such limitations may lead to incorrect results or failure to retrieve critical information, which can negatively impact decision-making in disaster situations.
+Reference Labeling Cohen’s $\kappa$
+Annotator 1 0.695
+Annotator 2 0.681
+Annotator 3 0.734
+Rounded average 0.690
+Majority vote (min in tie) 0.803
+Majority vote (max in tie) 0.715
+Fleiss’ $\kappa$ (inter-annotator) 0.777
+Table 6: Agreement between LLM-based relevance labels and human annotations on the in-domain sample (96 query–passage pairs).
 Information Need Generation Stage
@@
 System prompt: You are a search quality rater evaluating overall relevance.
-Given a query, passage, and sub-scores, provide a final score (0–3): 3 = Perfectly relevant, 2 = Highly relevant, 1 = Related, 0 = Irrelevant.
+Given a query, passage, and sub‐scores, provide a final score (0–3): 3 = Perfectly relevant, 2 = Highly relevant, 1 = Related, 0 = Irrelevant.
 Your output MUST be JSON: {"final_relevance_score": <0-3>}
-User prompt: Please rate how relevant the passage is based on the given sub-scores.Query: {query}, Passage: {passage}, Exactness: LLM score, Coverage: LLM score, Topicality: LLM score, Contextual Fit: LLM score.
+User prompt: Please rate how relevant the passage is based on the given sub‐scores.Query: {query}, Passage: {passage}, Exactness: LLM score, Coverage: LLM score, Topicality: LLM score, Contextual Fit: LLM score.
 Output your final rating as JSON: {"final_relevance_score": <integer_score_0_to_3>}
 Table 13: LLM relevance judgment prompt templates for QA, QAdoc, and Twitter-related search tasks.
-[1] Zero-shot Direct Scoring
+[1] Zero‐shot Direct Scoring
 System prompt: You are a search quality rater evaluating evidence for fact-checking and outputting JSON.
-User prompt: Given a claim and a passage, score relevance (003333): 00 = Irrelevant, $1$ = Related (no help), $2$ = Relevant (unclear/mixed), $3$ = Direct support/refutation.
+User prompt: Given a claim and a passage, score relevance (033): 0 = Irrelevant, $1$ = Related (no help), $2$ = Relevant (unclear/mixed), $3$ = Direct support/refutation.
 Important: $1$ if related but no help, $2$ if important info + noise, $3$ if clearly supports/refutes.
 Claim: {Query}, Passage: {Passage}. Consider intent, support/refutation ($M$), trustworthiness ($T$), then decide final score ($O$). Output MUST be JSON: {"final_score": <0_to_3>}
-[2] Chain-of-Thought Decomposed Reasoning
+[2] Chain‐of‐Thought Decomposed Reasoning
 PHASE 1 – Answer Presence Prediction:
@@
 Claim: {Query}, Passage: {Passage}. Output JSON: {"has_answer": "Yes"} or {"has_answer": "No"}
-PHASE 2 – Fine-grained Criterion Scoring:
+PHASE 2 – Fine‐grained Criterion Scoring:
 System prompt: Please assess how well the provided passage serves as evidence for evaluating the claim.
-Use the following scoring scale (003333): 00 = Not relevant at all / No information provided.
-$1$ = Marginally relevant / Partially addresses the criterion.2222 = Fairly relevant / Adequately addresses the criterion.
+Use the following scoring scale (033): 0 = Not relevant at all / No information provided.
+$1$ = Marginally relevant / Partially addresses the criterion.22 = Fairly relevant / Adequately addresses the criterion.
 $3$ = Highly relevant / Fully satisfies the criterion.
@@
 System prompt when Phase 1 = "Yes": You are a rater evaluating evidence for fact-checking.
-Score ($2$ or $3$): $2$ = Highly relevant (some support/refutation), $3$ = Perfectly relevant (direct support/refutation).
+Score ($2$ or $3$): 22 = Highly relevant (some support/refutation), $3$ = Perfectly relevant (direct support/refutation).
 Your output MUST be JSON: {"relevance_score": <2_or_3>} System prompt when Phase 1 = "No": You are a rater evaluating evidence for fact-checking.
-Score (00 or $1$): 00 = Irrelevant, $1$ = Related (no support/refutation).
+Score (0 or $1$): 0 = Irrelevant, $1$ = Related (no support/refutation).
 Your output MUST be JSON: {"relevance_score": <0_or_1>}
@@
 User prompt when Phase 1 = "No":
-Passage irrelevant for evidence. Rate its relevance (00 or $1$). Claim: {Query}, Passage: {Passage}. Output JSON: {"relevance_score": <0_or_1>}
-[3] Multi-Dimensional Attribute Scoring
-PHASE 1 – Sub-criterion Scoring:
+Passage irrelevant for evidence. Rate its relevance (0 or $1$). Claim: {Query}, Passage: {Passage}. Output JSON: {"relevance_score": <0_or_1>}
+[3] Multi‐Dimensional Attribute Scoring
+PHASE 1 – Sub‐criterion Scoring:
 System prompt: Please assess how well the provided passage serves as evidence for evaluating the claim according to the following criteria.
-Use a score (003333): 00 = Not relevant at all / No information provided.
+Use a score (033): 0 = Not relevant at all / No information provided.
 $1$ = Marginally relevant / Partially addresses the criterion.
@@
 System prompt: You are a search quality rater evaluating evidence relevance for fact-checking.
-Given a claim, passage, and sub-scores, provide a final score (003333): $3$ = Perfectly relevant (direct support/refutation), $2$ = Highly relevant (helps verification), $1$ = Related (topic match, no verification aid), 00 = Irrelevant.
+Given a claim, passage, and sub‐scores, provide a final score (033): 33 = Perfectly relevant (direct support/refutation), $2$ = Highly relevant (helps verification), $1$ = Related (topic match, no verification aid), 0 = Irrelevant.
 Your output MUST be JSON: {"final_relevance_score": <0_to_3>}
@@
 Table 14: LLM relevance judgment prompt templates for FC-related search tasks.
-[1] Zero-shot Direct Scoring
+[1] Zero‐shot Direct Scoring
 System prompt: You are a rater evaluating entailment. Output only the final score in JSON.
-User prompt: Given a premise and hypothesis, score entailment (003333): 00 = Not entailed/Contradicted, $1$ = Related not entailed, $2$ = Mostly entailed, $3$ = Perfectly entailed.
+User prompt: Given a premise and hypothesis, score entailment (033): 0 = Not entailed/Contradicted, $1$ = Related not entailed, $2$ = Mostly entailed, $3$ = Perfectly entailed.
 Important: $1$ if related but not inferable, $2$ if captures important implied content but not fully, $3$ if clearly/fully supported.
 Premise: {Query}, Hypothesis: {Passage}. Consider premise implications, logical following ($E$), info gaps, then decide final score ($O$). Output MUST be JSON: {"final_score": <0_to_3>}
-[2] Chain-of-Thought Decomposed Reasoning
+[2] Chain‐of‐Thought Decomposed Reasoning
 PHASE 1 – Answer Presence Prediction:
@@
 Premise: {Query}, Hypothesis: {Passage}.Output JSON: {"has_answer": "Yes"} or {"has_answer": "No"}
-PHASE 2 – Fine-grained Criterion Scoring:
+PHASE 2 – Fine‐grained Criterion Scoring:
 System prompt: Please assess how well the given hypothesis meets the {criterion_name} criterion in relation to the premise.
@@
 System prompt when Phase 1 = "Yes": You are a rater evaluating entailment.
-Provide a final score ($2$ or $3$): $2$ = Mostly entailed, $3$ = Perfectly entailed.
+Provide a final score ($2$ or $3$): 22 = Mostly entailed, $3$ = Perfectly entailed.
 Your output MUST be JSON: {"relevance_score": <2_or_3>}
 System prompt when Phase 1 = "No": You are a rater evaluating entailment.
-Provide a final score (00 or $1$): 00 = Not entailed/Contradicted, $1$ = Related but not entailed.
+Provide a final score (0 or $1$): 0 = Not entailed/Contradicted, $1$ = Related but not entailed.
 Your output MUST be JSON: {"relevance_score": <0_or_1>}
 User prompt when Phase 1 = "Yes": Hypothesis is entailed. Rate how well ($2$ or $3$). Premise: {Query}, Hypothesis: {Passage}. Output JSON: {"relevance_score": <2_or_3>}
-User prompt when Phase 1 = "No": Hypothesis is not entailed. Rate how (00 or $1$). Premise: {Query}, Hypothesis: {Passage}. Output JSON: {"relevance_score": <0_or_1>}
-[3] Multi-Dimensional Attribute Scoring
-PHASE 1 – Sub-criterion Scoring:
+User prompt when Phase 1 = "No": Hypothesis is not entailed. Rate how (0 or $1$). Premise: {Query}, Hypothesis: {Passage}. Output JSON: {"relevance_score": <0_or_1>}
+[3] Multi‐Dimensional Attribute Scoring
+PHASE 1 – Sub‐criterion Scoring:
 System prompt: Please assess how well the hypothesis is entailed by the premise according to the following criteria.
-Use a score (003333): 00 = Not relevant at all / No information provided.
+Use a score (033): 0 = Not relevant at all / No information provided.
 $1$ = Marginally relevant / Partially addresses the criterion.
@@
 System prompt: You are a search quality rater evaluating entailment.
-Given a premise, hypothesis, and sub-scores, provide a final score (003333): $3$ = Perfectly entailed, $2$ = Mostly entailed, $1$ = Related but not entailed, 00 = Not entailed/Contradicted.
+Given a premise, hypothesis, and sub‐scores, provide a final score (033): 33 = Perfectly entailed, $2$ = Mostly entailed, $1$ = Related but not entailed, 0 = Not entailed/Contradicted.
 Your output MUST be JSON: {"final_relevance_score": <0_to_3>}
@@
 Table 15: LLM relevance judgment prompt templates for NLI-related search tasks.
-[1] Zero-shot Direct Scoring for STS
-System prompt: You are a semantic-similarity rater. Output JSON with a score (005555) where 00 = unrelated, $5$ = semantically equivalent.
-User prompt: Rate the semantic similarity 005555: 00 = unrelated $\mid$ $1$ = slight $\mid$ $2$ = partial $\mid$ $3$ = moderate $\mid$ $4$ = high $\mid$ $5$ = equivalent
+[1] Zero‐shot Direct Scoring for STS
+System prompt: You are a semantic‐similarity rater. Output JSON with a score (055) where 0 = unrelated, $5$ = semantically equivalent.
+User prompt: Rate the semantic similarity 055: 0 = unrelated $\mid$ 11 = slight $\mid$ 22 = partial $\mid$ 33 = moderate $\mid$ 44 = high $\mid$ 55 = equivalent
 Sentence A: {input1} Sentence B: {input2}
@@
 snowflake-arctic-embed-m 109M Medium e5-unsupervised-base 764 13 encoder
-granite-embedding-125m-english 125M Medium RoBERTa 768 14 encoder
-bge-large-en-v1.5 335M Large 1024 15 encoder
-mxbai-embed-large-v1 335M Large 1024 16 encoder
-snowflake-arctic-embed-s 33M Small e5-unsupervised-small 384 17 encoder
-bge-base-en-v1.5 109M Medium 768 18 encoder
-bge-small-en-v1.5 33M Small 384 19 encoder
-multilingual-e5-large-instruct 560M Large xlm-roberta-large 1024 20 encoder
-thenlper-gte-base 109M Medium EBRT-base 768 21 encoder
-multilingual-e5-large 560M Large xlm-roberta-large 1024 22 encoder
-gte-Qwen2-1.5B-instruct 1.5B XL Qwen2-1.5B 1536 23 decoder
-e5-base-v2 109M Medium bert-large-uncased 1024 24 encoder
-e5-large-v2 335M Large bert-base-uncased 768 25 encoder
-e5-small-v2 33M Small MiniLM 384 26 encoder
-all-mpnet-base-v2 109M Medium microsoft/mpnet-base 768 27 encoder
-gte-base-en-v1.5 137M Medium EBRT-base 768 28 encoder
+snowflake-arctic-embed-m-long 137M Medium nomic-embed-text-v1-unsupervised 768 14 encoder
+granite-embedding-125m-english 125M Medium RoBERTa 768 15 encoder
+bge-large-en-v1.5 335M Large 1024 16 encoder
+mxbai-embed-large-v1 335M Large 1024 17 encoder
+snowflake-arctic-embed-s 33M Small e5-unsupervised-small 384 18 encoder
+bge-base-en-v1.5 109M Medium 768 19 encoder
+bge-small-en-v1.5 33M Small 384 20 encoder
+multilingual-e5-large-instruct 560M Large xlm-roberta-large 1024 21 encoder
+thenlper-gte-base 109M Medium EBRT-base 768 22 encoder
+multilingual-e5-large 560M Large xlm-roberta-large 1024 23 encoder
+thenlper-gte-small 33M Small MiniLM-L12-H384 384 24 encoder
+gte-Qwen2-1.5B-instruct 1.5B XL Qwen2-1.5B 1536 25 decoder
+e5-base-v2 109M Medium bert-large-uncased 1024 26 encoder
+e5-large-v2 335M Large bert-base-uncased 768 27 encoder
+e5-small-v2 33M Small MiniLM 384 28 encoder
+gte-base-en-v1.5 137M Medium EBRT-base 768 29 encoder
 gte-large-en-v1.5 434M Large EBRT-large 1024 encoder
-llmrails-ember-v1 335M Large 1024 encoder
 Table 20: Information of all evaluated models. “–” means no publicly available information is available.
@@
 inf-retriever-v1 https://huggingface.co/infly/inf-retriever-v1 apache-2.0
-NV-Embed-v2 https://huggingface.co/nvidia/NV-Embed-v2 cc-by-nc-4.0
+NV-Embed-v2 https://huggingface.co/nvidia/NV-Embed-v2 $c$-by-nc-4.0
 inf-retriever-v1-1.5b https://huggingface.co/infly/inf-retriever-v1-1.5b apache-2.0
-Linq-Embed-Mistral https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral cc-by-nc-4.0
-NV-Embed-v1 https://huggingface.co/nvidia/NV-Embed-v1 cc-by-nc-4.0
-SFR-Embedding-Mistral https://huggingface.co/Salesforce/SFR-Embedding-Mistral cc-by-nc-4.0
+Linq-Embed-Mistral https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral $c$-by-nc-4.0
+NV-Embed-v1 https://huggingface.co/nvidia/NV-Embed-v1 $c$-by-nc-4.0
+SFR-Embedding-Mistral https://huggingface.co/Salesforce/SFR-Embedding-Mistral $c$-by-nc-4.0
 snowflake-arctic-embed-l https://huggingface.co/Snowflake/snowflake-arctic-embed-l apache-2.0
@@
 snowflake-arctic-embed-m https://huggingface.co/Snowflake/snowflake-arctic-embed-m apache-2.0
+snowflake-arctic-embed-m https://huggingface.co/Snowflake/snowflake-arctic-embed-m-long apache-2.0
 granite-embedding-125m-english https://huggingface.co/ibm-granite/granite-embedding-125m-english mit
@@
 multilingual-e5-large https://huggingface.co/intfloat/multilingual-e5-large-instructt apache-2.0
+thenlper-gte-small https://huggingface.co/thenlper/gte-small mit
 gte-Qwen2-1.5B-instruct https://huggingface.co/Alibaba-NLP/gte-Qwen2-1.5B-instruct mit
@@
 e5-small-v2 https://huggingface.co/intfloat/e5-small-v2 apache-2.0
-all-mpnet-base-v2 https://huggingface.co/sentence-transformers/all-mpnet-base-v2 apache-2.0
 gte-base-en-v1.5 https://huggingface.co/Alibaba-NLP/gte-base-en-v1.5 apache-2.0
 gte-large-en-v1.5 https://huggingface.co/Alibaba-NLP/gte-large-en-v1.5 mit
-llmrails-ember-v1 https://huggingface.co/llmrails/ember-v1 mit
 Table 21: HuggingFace model links and licenses for all evaluated models.
@@
 Alibaba-NLP-gte-Qwen2-1.5B-instruct
-Biological 12.93 19.75 19.16 25.18 18.82 34.19 21.67
-Chemical 14.74 20.20 19.57 25.92 15.82 26.73 20.50
-Environmental 14.03 26.03 23.66 21.48 19.75 34.39 23.23
-Extraterrestrial 13.13 21.89 15.62 27.12 17.38 30.09 20.87
-Geohazard 14.68 17.35 18.51 21.79 15.51 34.44 20.38
-Meteorological&hydrological 12.63 19.33 23.63 22.10 19.47 30.22 21.23
-Societal 14.51 29.00 17.56 24.04 18.59 22.57 21.05
-Technological 15.20 24.16 19.15 23.58 18.63 37.00 22.95
-Avg. 13.98 22.21 19.61 23.90 18.00 31.20 21.48
+Biological 70.58 60.91 68.01 61.43 55.36 74.59 65.15
+Chemical 70.02 58.58 64.21 64.40 55.48 73.43 64.35
+Environmental 69.88 56.90 65.41 64.73 59.01 73.74 64.94
+Extraterrestrial 67.56 56.18 61.33 62.44 53.14 74.72 62.56
+Geohazard 70.30 58.34 63.81 59.60 54.98 71.91 63.16
+Meteorological&hydrological 70.45 59.65 66.28 59.56 54.95 72.93 63.97
+Societal 71.41 61.67 67.07 64.68 58.40 73.51 66.12
+Technological 69.44 61.46 65.57 65.90 54.54 74.01 65.16
+Avg. 69.96 59.21 65.21 62.84 55.73 73.61 64.43
 Alibaba-NLP-gte-Qwen2-7B-instruct
-Biological 62.58 37.35 43.67 27.64 45.76 71.38 48.06
-Chemical 64.42 38.19 41.06 31.31 46.84 70.43 48.71
-Environmental 62.58 38.36 41.83 29.77 48.23 71.06 48.64
-Extraterrestrial 58.55 34.42 37.55 31.45 46.54 73.40 46.99
-Geohazard 61.12 36.31 40.13 28.31 47.12 68.06 46.84
-Meteorological&hydrological 62.22 39.15 42.08 26.14 45.62 70.57 47.63
-Societal 66.77 47.63 43.10 31.95 51.06 72.05 52.09
-Technological 61.51 41.51 41.80 33.27 47.08 70.86 49.34
-Avg. 62.47 39.12 41.40 29.98 47.28 70.98 48.54
+Biological 70.42 47.68 64.56 32.33 53.79 75.93 57.45
+Chemical 71.77 47.98 62.94 35.13 52.81 75.01 57.61
+Environmental 69.36 44.00 63.04 30.11 56.15 75.33 56.33
+Extraterrestrial 67.12 44.91 59.87 32.20 50.51 74.59 54.87
+Geohazard 71.17 47.91 63.11 30.85 53.78 74.34 56.86
+Meteorological&hydrological 70.30 48.08 64.94 27.68 52.61 73.10 56.12
+Societal 71.81 50.68 64.35 32.02 57.89 74.82 58.59
+Technological 70.45 49.94 63.06 34.64 53.53 75.80 57.90
+Avg. 70.30 47.65 63.24 31.87 53.88 74.86 56.97
 Alibaba-NLP-gte-base-en-v1.5
-Biological 60.86 54.84 46.25 53.77 42.53 70.95 54.87
-Chemical 60.99 55.23 43.96 53.63 37.30 70.23 53.56
-Environmental 62.02 55.77 47.11 51.32 43.30 71.72 55.21
-Extraterrestrial 59.48 53.38 49.06 53.30 37.65 72.35 54.20
-Geohazard 60.10 54.26 45.96 49.82 36.83 70.45 52.90
-Meteorological&hydrological 59.98 58.31 47.26 51.44 39.13 70.93 54.51
-Societal 61.99 59.05 48.04 53.11 42.88 71.83 56.15
-Technological 59.83 57.66 48.85 53.55 39.39 71.68 55.16
-Avg. 60.66 56.06 47.06 52.49 39.88 71.27 54.57
+Biological 60.71 54.57 45.27 53.77 42.13 70.02 54.41
+Chemical 60.83 54.93 43.47 53.53 37.44 69.37 53.26
+Environmental 61.88 55.46 46.43 51.08 43.30 71.11 54.88
+Extraterrestrial 59.20 53.20 48.45 53.15 37.86 71.54 53.90
+Geohazard 60.00 54.10 45.55 49.61 37.01 69.57 52.64
+Meteorological&hydrological 59.74 58.13 46.70 51.29 39.06 69.79 54.12
+Societal 61.78 58.93 47.54 52.86 42.46 70.98 55.76
+Technological 59.54 57.48 48.11 53.45 39.55 70.89 54.84
+Avg. 60.46 55.85 46.44 52.34 39.85 70.41 54.23
 Alibaba-NLP-gte-large-en-v1.5
-Biological 68.22 59.69 43.28 56.01 36.94 67.79 55.32
-Chemical 70.28 59.30 38.93 56.44 33.85 68.50 54.55
-Environmental 67.62 57.23 41.34 52.00 38.74 66.61 53.92
-Extraterrestrial 65.28 56.57 42.45 51.24 32.11 68.85 52.75
-Geohazard 65.65 52.61 35.16 49.45 30.81 64.75 49.74
-Meteorological&hydrological 67.20 59.98 38.81 51.14 33.15 66.41 52.78
-Societal 69.23 62.26 41.72 53.73 38.01 68.27 55.54
-Technological 68.15 60.74 40.77 54.70 34.90 67.23 54.41
-Avg. 67.70 58.55 40.31 53.09 34.81 67.30 53.63
+Biological 68.08 59.51 42.33 55.85 37.09 66.90 54.96
+Chemical 70.12 59.08 38.44 56.34 33.66 67.67 54.22
+Environmental 67.34 56.92 40.80 51.83 39.08 66.24 53.70
+Extraterrestrial 65.06 56.49 41.82 51.08 32.19 68.13 52.46
+Geohazard 65.43 52.43 34.88 49.19 30.74 63.90 49.43
+Meteorological&hydrological 66.95 59.82 38.31 50.95 32.76 65.32 52.35
+Societal 68.89 62.11 41.13 53.47 37.61 67.46 55.11
+Technological 67.83 60.59 39.96 54.51 35.22 66.47 54.10
+Avg. 67.46 58.37 39.71 52.90 34.79 66.51 53.29
 BAAI-bge-base-en-v1.5
-Biological 51.58 53.85 50.86 58.41 37.78 65.45 52.99
-Chemical 47.97 50.77 48.75 58.64 34.71 67.10 51.32
-Environmental 50.63 51.44 47.75 60.93 41.00 63.61 52.56
-Extraterrestrial 50.50 57.93 45.21 59.49 33.18 65.59 51.98
-Geohazard 47.81 52.76 44.28 56.91 33.09 62.52 49.56
-Meteorological&hydrological 49.27 53.06 47.10 59.05 34.13 62.09 50.78
-Societal 50.00 59.17 45.32 55.97 36.82 62.81 51.68
-Technological 49.56 52.77 46.99 58.39 33.90 66.24 51.31
-Avg. 49.67 53.97 47.03 58.47 35.58 64.43 51.52
+Biological 53.05 53.49 50.94 60.72 42.05 69.93 55.03
+Chemical 50.87 49.40 49.24 59.74 40.53 70.65 53.40
+Environmental 52.48 51.20 46.84 63.06 46.04 68.80 54.74
+Extraterrestrial 52.19 56.57 45.08 59.60 39.80 69.14 53.73
+Geohazard 49.40 50.32 43.49 58.21 39.90 66.28 51.27
+Meteorological&hydrological 51.49 51.47 46.42 60.84 39.81 66.60 52.77
+Societal 52.35 58.35 45.70 59.23 43.55 67.96 54.52
+Technological 51.40 52.30 46.54 59.64 39.59 69.15 53.10
+Avg. 51.65 52.89 46.78 60.13 41.41 68.56 53.57
 BAAI-bge-large-en-v1.5
-Biological 56.35 55.10 36.71 55.50 35.62 66.83 51.02
-Chemical 56.24 51.49 33.57 56.73 34.65 66.02 49.78
-Environmental 57.20 51.53 33.76 54.54 38.85 65.74 50.27
-Extraterrestrial 58.77 57.02 33.03 58.30 32.68 66.38 51.03
-Geohazard 55.15 53.62 30.65 53.17 35.16 63.51 48.54
-Meteorological&hydrological 56.45 54.24 31.36 53.73 33.20 62.50 48.58
-Societal 58.61 58.52 32.33 54.25 35.96 63.75 50.57
-Technological 57.39 56.25 30.28 55.56 35.44 66.37 50.21
-Avg. 57.02 54.72 32.71 55.22 35.19 65.14 50.00
+Biological 56.20 54.80 36.10 55.29 35.67 66.15 50.70
+Chemical 56.24 51.57 33.14 56.65 34.44 65.21 49.54
+Environmental 57.11 51.28 33.24 54.35 38.65 65.29 49.99
+Extraterrestrial 58.73 56.79 32.68 58.04 33.39 65.66 50.88
+Geohazard 55.18 53.50 30.47 52.89 35.70 62.73 48.41
+Meteorological&hydrological 56.09 54.19 31.05 53.69 33.13 61.63 48.30
+Societal 58.30 58.31 32.08 54.01 35.92 63.03 50.28
+Technological 57.18 56.05 29.83 55.32 35.08 65.75 49.87
+Avg. 56.88 54.56 32.32 55.03 35.25 64.43 49.75
 Table 22: Performance of the first six evaluated models under six search intents and eight event types under the exact search setting. Part I
@@
 BAAI-bge-small-en-v1.5
-Biological 47.81 47.68 31.38 53.01 28.93 60.77 44.93
-Chemical 45.25 42.32 30.68 55.45 27.44 57.61 43.12
-Environmental 51.27 45.90 28.36 52.48 31.09 53.90 43.83
-Extraterrestrial 48.32 45.50 26.64 53.35 25.57 58.43 42.97
-Geohazard 46.74 42.91 25.79 50.88 25.91 55.30 41.25
-Meteorological&hydrological 46.89 46.96 26.52 52.61 28.11 55.46 42.76
-Societal 47.90 47.18 26.05 45.50 25.78 54.76 41.20
-Technological 44.82 45.61 27.12 53.74 24.75 58.68 42.45
-Avg. 47.38 45.51 27.82 52.13 27.20 56.86 42.82
+Biological 58.08 52.41 28.12 55.48 35.55 66.01 49.28
+Chemical 56.12 48.37 27.78 57.54 32.94 64.88 47.94
+Environmental 60.67 51.59 24.29 56.89 37.56 64.23 49.21
+Extraterrestrial 56.59 51.14 23.53 56.96 32.45 65.57 47.71
+Geohazard 54.58 48.79 22.84 53.02 30.27 63.40 45.48
+Meteorological&hydrological 56.82 51.60 24.47 54.81 32.21 62.84 47.13
+Societal 58.51 54.46 25.09 51.98 31.96 63.81 47.63
+Technological 53.59 51.55 25.36 55.68 30.71 64.93 46.97
+Avg. 56.87 51.24 25.19 55.30 32.95 64.46 47.67
 Linq-AI-Research-Linq-Embed-Mistral
-Biological 35.16 31.88 27.70 43.16 28.81 50.09 36.13
-Chemical 39.05 28.91 27.52 44.70 20.32 39.50 33.33
-Environmental 42.55 33.30 30.97 36.88 30.97 50.49 37.53
-Extraterrestrial 33.98 33.26 25.64 44.49 28.29 42.68 34.72
-Geohazard 35.77 28.05 23.63 33.08 27.89 47.75 32.70
-Meteorological&hydrological 33.17 28.16 28.04 32.61 28.43 44.46 32.48
-Societal 35.09 36.96 23.60 38.97 30.61 39.84 34.18
-Technological 30.12 30.78 25.34 41.78 23.06 50.72 33.63
-Avg. 35.61 31.41 26.55 39.46 27.30 45.69 34.34
+Biological 75.63 71.17 66.30 70.33 52.99 71.98 68.07
+Chemical 74.96 69.32 63.35 72.40 50.04 70.41 66.75
+Environmental 74.33 68.53 65.57 72.02 55.38 71.09 67.82
+Extraterrestrial 73.32 71.07 62.56 72.36 51.42 73.09 67.30
+Geohazard 73.79 68.30 62.62 68.15 51.33 68.24 65.40
+Meteorological&hydrological 74.41 70.76 63.94 70.04 52.57 71.04 67.13
+Societal 75.17 74.31 65.18 69.84 54.54 72.31 68.56
+Technological 73.53 70.53 64.22 71.01 52.24 72.65 67.36
+Avg. 74.40 70.50 64.22 70.77 52.56 71.35 67.30
 Salesforce-SFR-Embedding-Mistral
-Biological 72.89 67.94 72.41 69.18 51.66 74.32 68.07
-Chemical 72.03 66.49 69.62 72.05 49.51 72.65 67.06
-Environmental 72.25 66.24 72.43 73.16 52.99 73.20 68.38
-Extraterrestrial 70.01 68.71 67.85 70.71 47.73 75.48 66.75
-Geohazard 70.77 65.41 69.06 69.21 49.01 70.81 65.71
-Meteorological&hydrological 72.44 67.29 70.11 70.49 50.84 73.12 67.38
-Societal 72.78 70.13 71.38 69.38 53.92 74.26 68.64
-Technological 70.89 68.42 70.48 70.80 51.20 75.08 67.81
-Avg. 71.76 67.58 70.42 70.62 50.86 73.61 67.47
+Biological 72.68 67.70 71.20 69.09 52.40 73.38 67.74
+Chemical 71.82 66.39 68.83 71.86 49.62 71.69 66.70
+Environmental 72.08 65.97 71.56 72.98 53.07 72.53 68.03
+Extraterrestrial 69.81 68.26 66.93 70.59 48.33 74.66 66.43
+Geohazard 70.61 65.05 68.54 68.63 49.72 69.86 65.40
+Meteorological&hydrological 72.13 67.12 69.66 70.34 51.06 71.88 67.03
+Societal 72.47 69.89 70.67 69.04 53.50 73.32 68.15
+Technological 70.40 68.35 69.58 70.57 50.94 74.36 67.37
+Avg. 71.50 67.34 69.62 70.39 51.08 72.71 67.11
 Snowflake-snowflake-arctic-embed-l
-Biological 41.34 30.98 16.73 33.28 35.21 54.92 35.41
-Chemical 40.77 30.26 16.39 34.55 33.24 56.39 35.27
-Environmental 43.56 34.71 17.68 34.42 40.28 56.81 37.91
-Extraterrestrial 41.09 27.96 14.48 31.93 32.31 58.93 34.45
-Geohazard 36.34 28.20 12.92 33.83 31.08 53.90 32.71
-Meteorological&hydrological 37.75 25.59 14.99 33.70 34.90 57.75 34.11
-Societal 42.99 35.02 16.49 28.39 36.37 57.69 36.16
-Technological 42.70 30.53 12.85 31.51 33.57 58.20 34.89
-Avg. 40.82 30.41 15.32 32.70 34.62 56.82 35.11
+Biological 41.19 30.92 16.44 33.21 34.85 54.28 35.15
+Chemical 40.74 30.23 16.05 34.55 33.12 55.60 35.05
+Environmental 43.49 34.59 17.50 34.29 40.35 56.38 37.77
+Extraterrestrial 41.20 28.02 14.27 31.82 32.23 58.17 34.28
+Geohazard 36.39 28.06 12.86 33.70 31.00 53.15 32.53
+Meteorological&hydrological 37.65 25.44 14.77 33.56 34.65 56.80 33.81
+Societal 42.77 34.94 16.33 28.26 36.31 56.89 35.92
+Technological 42.39 30.46 12.67 31.41 33.02 57.57 34.59
+Avg. 40.73 30.33 15.11 32.60 34.44 56.11 34.89
 Snowflake-snowflake-arctic-embed-l-v2.0
-Biological 53.33 58.87 38.42 60.63 41.22 62.01 52.41
-Chemical 56.40 57.57 40.62 61.11 39.01 63.87 53.10
-Environmental 57.52 59.91 45.04 62.53 45.87 63.73 55.77
-Extraterrestrial 54.74 56.70 36.38 60.14 37.86 62.82 51.44
-Geohazard 54.21 57.26 34.76 59.82 39.80 60.19 51.01
-Meteorological&hydrological 51.69 59.76 36.28 61.42 40.67 62.49 52.05
-Societal 57.78 63.46 41.87 57.77 44.56 64.80 55.04
-Technological 57.02 61.94 36.67 59.97 40.49 67.14 53.87
-Avg. 55.34 59.43 38.75 60.42 41.19 63.38 53.09
+Biological 53.19 58.69 37.59 60.60 41.47 61.42 52.16
+Chemical 56.34 57.58 40.17 60.94 39.15 63.01 52.86
+Environmental 57.44 59.63 44.30 62.25 46.12 63.24 55.50
+Extraterrestrial 54.62 56.50 35.98 60.03 38.19 62.14 51.24
+Geohazard 54.13 57.02 34.45 59.46 39.61 59.40 50.68
+Meteorological&hydrological 51.56 59.75 35.89 61.26 40.85 61.52 51.80
+Societal 57.58 63.32 41.46 57.55 44.13 64.00 54.67
+Technological 56.77 61.82 36.28 59.77 40.34 66.39 53.56
+Avg. 55.20 59.29 38.26 60.23 41.23 62.64 52.81
 Snowflake-snowflake-arctic-embed-m
-Biological 31.68 13.54 8.09 38.04 39.39 54.55 30.88
-Chemical 33.35 14.43 9.76 35.38 36.26 58.17 31.23
-Environmental 35.53 14.87 9.18 35.83 43.61 57.19 32.70
-Extraterrestrial 33.40 14.31 8.29 35.18 39.67 58.59 31.57
-Geohazard 31.14 14.32 6.15 35.78 36.75 54.06 29.70
-Meteorological&hydrological 30.96 14.49 9.29 35.75 39.04 56.34 30.98
-Societal 35.17 14.34 10.72 31.87 40.57 57.82 31.75
-Technological 35.55 13.70 7.63 34.58 36.14 58.32 30.99
-Avg. 33.35 14.25 8.64 35.30 38.93 56.88 31.23
+Biological 31.64 13.53 7.84 37.99 39.26 54.05 30.72
+Chemical 33.30 14.33 9.69 35.25 36.09 57.36 31.00
+Environmental 35.37 14.95 9.07 35.77 43.12 56.79 32.51
+Extraterrestrial 33.43 14.24 8.24 34.96 39.61 57.93 31.40
+Geohazard 31.11 14.33 6.01 35.51 36.50 53.24 29.45
+Meteorological&hydrological 30.88 14.45 9.19 35.66 38.65 55.44 30.71
+Societal 35.04 14.23 10.67 31.79 40.81 57.21 31.62
+Technological 35.36 13.69 7.44 34.34 35.96 57.68 30.74
+Avg. 33.26 14.22 8.52 35.16 38.75 56.21 31.02
 Table 23: Performance of evaluated models under six search intents and eight event types under the exact search setting. Part II
@@
 Snowflake-snowflake-arctic-embed-m-v1.5
-Biological 20.41 30.10 17.56 48.74 42.93 63.69 37.24
-Chemical 23.39 29.15 20.20 48.55 41.74 65.45 38.08
-Environmental 28.64 33.80 21.74 51.30 49.00 65.53 41.67
-Extraterrestrial 25.43 24.07 17.35 46.65 41.79 66.04 36.89
-Geohazard 26.06 30.31 17.01 46.63 42.02 63.49 37.59
-Meteorological&hydrological 22.61 28.98 17.43 48.37 42.51 64.42 37.39
-Societal 29.55 35.48 17.74 47.59 42.61 64.06 39.51
-Technological 29.77 32.56 17.47 48.12 41.71 66.91 39.42
-Avg. 25.73 30.56 18.31 48.24 43.04 64.95 38.47
+Biological 20.43 30.04 17.22 48.66 43.01 63.04 37.07
+Chemical 23.35 29.02 19.92 48.41 41.47 64.66 37.80
+Environmental 28.52 33.59 21.38 51.31 48.85 65.02 41.45
+Extraterrestrial 25.42 24.03 17.19 46.50 41.84 65.34 36.72
+Geohazard 25.97 30.09 16.87 46.30 41.72 62.68 37.27
+Meteorological&hydrological 22.59 28.84 17.24 48.22 42.64 63.29 37.14
+Societal 29.41 35.38 17.63 47.40 42.67 63.35 39.31
+Technological 29.59 32.49 17.24 47.99 41.66 66.24 39.20
+Avg. 25.66 30.43 18.09 48.10 42.98 64.20 38.24
 Snowflake-snowflake-arctic-embed-m-v2.0
-Biological 60.58 62.53 48.09 57.58 42.68 64.84 56.05
-Chemical 61.22 61.01 47.12 59.78 40.78 64.72 55.77
-Environmental 63.96 62.03 49.99 60.78 47.75 65.53 58.34
-Extraterrestrial 60.83 61.17 46.86 58.68 39.05 65.74 55.39
-Geohazard 60.38 60.30 47.90 57.05 43.33 63.02 55.33
-Meteorological&hydrological 60.47 63.37 46.95 58.29 40.69 63.79 55.59
-Societal 62.72 65.79 47.63 54.61 43.17 66.73 56.78
-Technological 61.34 63.79 47.86 57.56 41.24 67.88 56.61
-Avg. 61.44 62.50 47.80 58.04 42.34 65.28 56.23
+Biological 60.41 62.36 47.15 57.49 42.82 64.21 55.74
+Chemical 61.13 60.93 46.65 59.61 40.91 64.03 55.54
+Environmental 63.92 61.70 49.28 60.57 47.70 65.08 58.04
+Extraterrestrial 60.68 60.86 46.24 58.47 39.51 65.04 55.13
+Geohazard 60.29 60.11 47.50 56.66 43.53 62.19 55.05
+Meteorological&hydrological 60.22 63.23 46.42 58.08 40.62 62.84 55.23
+Societal 62.52 65.63 47.14 54.42 43.23 65.96 56.48
+Technological 61.07 63.65 47.19 57.40 41.10 67.18 56.27
+Avg. 61.28 62.31 47.20 57.84 42.43 64.56 55.94
 Snowflake-snowflake-arctic-embed-s
-Biological 34.80 26.88 18.93 44.11 36.25 66.25 37.87
-Chemical 33.78 28.03 20.01 45.40 35.16 66.72 38.18
-Environmental 39.09 30.43 23.33 42.75 43.67 68.17 41.24
-Extraterrestrial 35.15 21.60 16.03 41.94 34.80 66.06 35.93
-Geohazard 35.78 25.98 16.63 41.53 36.37 63.31 36.60
-Meteorological&hydrological 34.13 28.33 16.74 43.00 38.08 66.33 37.77
-Societal 37.91 30.80 17.34 39.73 38.71 67.63 38.69
-Technological 37.77 27.63 16.34 41.68 35.15 68.29 37.81
-Avg. 36.05 27.46 18.17 42.52 37.27 66.59 38.01
+Biological 38.45 28.71 22.49 49.59 39.71 66.50 40.91
+Chemical 36.35 29.04 23.56 49.73 37.62 66.73 40.51
+Environmental 41.72 30.95 26.34 48.17 46.21 68.23 43.60
+Extraterrestrial 37.42 22.95 18.23 46.11 37.85 66.97 38.26
+Geohazard 37.63 26.73 19.34 45.56 38.60 64.43 38.71
+Meteorological&hydrological 37.03 29.92 21.05 47.02 39.78 66.11 40.15
+Societal 41.05 32.44 21.38 45.08 41.98 67.94 41.65
+Technological 39.89 29.84 19.03 47.15 38.40 68.66 40.50
+Avg. 38.69 28.82 21.43 47.30 40.02 66.95 40.54
 ibm-granite-granite-embedding-125m-english
-Biological 65.94 61.42 50.92 62.06 46.86 71.23 59.74
-Chemical 64.39 57.94 52.68 63.52 46.68 71.24 59.41
-Environmental 65.87 60.32 47.60 64.84 52.70 73.30 60.77
-Extraterrestrial 63.75 58.73 44.83 62.05 47.34 72.81 58.25
-Geohazard 63.08 59.78 44.29 60.60 46.28 70.89 57.49
-Meteorological&hydrological 65.83 61.04 45.60 62.71 45.98 71.02 58.70
-Societal 66.36 65.53 44.21 62.98 50.31 71.93 60.22
-Technological 63.95 63.52 47.03 63.49 48.21 73.06 59.88
-Avg. 64.90 61.04 47.14 62.78 48.04 71.94 59.31
+Biological 65.76 61.24 49.90 61.91 46.66 70.40 59.31
+Chemical 64.14 57.95 52.14 63.28 46.64 70.31 59.08
+Environmental 65.74 60.20 47.01 64.65 52.02 72.71 60.39
+Extraterrestrial 63.53 58.39 44.19 61.87 48.12 72.01 58.02
+Geohazard 62.93 59.57 43.92 60.19 46.92 69.93 57.25
+Meteorological&hydrological 65.40 60.87 45.27 62.50 46.00 69.76 58.30
+Societal 66.06 65.24 43.73 62.84 50.83 71.08 59.96
+Technological 63.50 63.37 46.26 63.20 47.71 72.31 59.39
+Avg. 64.63 60.85 46.55 62.56 48.11 71.06 58.96
 infly-inf-retriever-v1
-Biological 76.15 68.85 71.20 68.79 52.90 78.79 69.45
-Chemical 74.65 65.12 68.88 70.95 52.85 76.74 68.20
-Environmental 73.45 66.91 69.84 70.05 56.45 77.55 69.04
-Extraterrestrial 72.61 68.14 66.38 68.78 48.90 79.31 67.35
-Geohazard 74.22 65.10 68.63 67.46 51.55 75.86 67.14
-Meteorological&hydrological 74.21 69.31 70.86 68.37 52.17 76.76 68.61
-Societal 75.49 70.65 69.64 68.07 57.23 78.49 69.93
-Technological 73.03 68.48 69.23 68.84 52.73 78.12 68.41
-Avg. 74.23 67.82 69.33 68.91 53.10 77.70 68.52
+Biological 74.92 68.15 68.16 65.94 52.10 77.21 67.75
+Chemical 73.23 64.74 66.32 67.34 51.42 75.34 66.40
+Environmental 72.43 65.78 65.62 66.67 54.90 76.13 66.92
+Extraterrestrial 70.64 66.50 63.10 64.90 47.66 77.47 65.04
+Geohazard 73.01 64.51 66.10 64.32 51.30 73.88 65.52
+Meteorological&hydrological 72.60 68.24 68.24 65.71 50.53 74.64 66.66
+Societal 74.06 69.47 67.21 65.80 55.77 76.62 68.15
+Technological 71.80 67.95 66.18 65.39 52.48 76.69 66.75
+Avg. 72.84 66.92 66.37 65.76 52.02 76.00 66.65
 infly-inf-retriever-v1-1.5b
-Biological 72.23 65.55 67.42 65.01 53.59 76.99 66.80
-Chemical 71.18 63.09 65.04 67.03 51.93 75.97 65.71
-Environmental 70.06 63.56 65.57 68.21 57.26 76.07 66.79
-Extraterrestrial 67.37 61.83 63.32 65.93 51.12 77.02 64.43
-Geohazard 71.49 62.52 65.70 65.56 52.29 75.73 65.55
-Meteorological&hydrological 70.52 65.69 67.86 67.18 52.54 75.66 66.58
-Societal 71.90 66.62 66.65 66.51 57.99 76.24 67.65
-Technological 70.39 66.48 67.29 66.79 54.18 75.83 66.83
-Avg. 70.64 64.42 66.11 66.53 53.86 76.19 66.29
+Biological 71.44 65.78 63.63 64.76 54.11 74.52 65.71
+Chemical 69.49 63.79 61.76 65.78 51.92 73.97 64.45
+Environmental 68.35 62.30 63.38 67.12 56.62 74.42 65.36
+Extraterrestrial 66.65 61.14 60.58 64.80 51.64 75.18 63.33
+Geohazard 70.18 62.68 62.51 63.96 53.24 73.25 64.31
+Meteorological&hydrological 69.75 65.58 64.54 66.51 53.45 72.88 65.45
+Societal 70.68 66.83 64.26 65.48 58.05 74.21 66.59
+Technological 69.22 67.09 64.00 65.52 54.13 73.28 65.54
+Avg. 69.47 64.40 63.08 65.49 54.14 73.96 65.09
 Table 24: Performance of evaluated models under six search intents and eight event types under the exact search setting. Part III
@@
 intfloat-e5-base-v2
-Biological 67.85 62.93 58.90 62.08 47.03 74.82 62.27
-Chemical 66.55 59.44 58.58 63.27 45.05 74.71 61.27
-Environmental 67.45 63.11 60.69 64.54 49.66 74.99 63.41
-Extraterrestrial 63.48 63.63 55.96 63.26 43.63 75.01 60.83
-Geohazard 65.17 60.34 56.01 60.72 43.17 74.34 59.96
-Meteorological&hydrological 66.14 62.64 59.13 62.21 44.88 74.08 61.52
-Societal 64.92 67.05 59.00 60.68 46.63 74.02 62.05
-Technological 64.58 65.38 58.65 61.44 44.27 75.15 61.58
-Avg. 65.77 63.07 58.37 62.28 45.54 74.64 61.61
+Biological 67.77 62.64 57.97 62.07 46.84 73.89 61.86
+Chemical 66.40 59.51 57.92 63.12 45.14 73.75 60.98
+Environmental 67.34 62.93 59.94 64.35 49.39 74.33 63.05
+Extraterrestrial 63.25 63.36 55.28 63.11 44.20 74.21 60.57
+Geohazard 65.05 60.12 55.61 60.40 43.29 73.40 59.64
+Meteorological&hydrological 65.83 62.57 58.66 62.12 44.30 72.80 61.04
+Societal 64.62 66.76 58.51 60.48 46.61 73.13 61.68
+Technological 64.06 65.39 58.19 61.24 44.36 74.32 61.26
+Avg. 65.54 62.91 57.76 62.11 45.52 73.73 61.26
 intfloat-e5-large-v2
-Biological 59.31 63.11 56.68 62.40 51.02 75.08 61.27
-Chemical 59.77 59.78 54.60 63.05 49.49 75.10 60.30
-Environmental 62.61 63.25 57.49 64.79 54.51 75.21 62.98
-Extraterrestrial 58.42 63.93 54.57 61.76 48.29 75.67 60.44
-Geohazard 59.25 62.26 54.94 60.25 50.44 74.02 60.19
-Meteorological&hydrological 60.07 63.55 57.59 62.98 49.77 75.12 61.51
-Societal 60.04 66.56 56.92 60.85 53.09 74.75 62.04
-Technological 61.74 64.95 56.80 62.20 51.07 74.97 61.95
-Avg. 60.15 63.42 56.20 62.29 50.96 74.99 61.34
+Biological 59.28 62.94 55.52 62.28 50.83 74.14 60.83
+Chemical 59.63 59.82 53.97 62.90 49.44 74.20 59.99
+Environmental 62.62 62.69 56.69 64.54 54.63 74.58 62.62
+Extraterrestrial 58.19 63.70 53.88 61.32 49.10 74.81 60.17
+Geohazard 59.19 62.11 54.49 59.83 50.74 73.16 59.92
+Meteorological&hydrological 59.85 63.30 56.85 62.79 49.28 73.84 60.98
+Societal 59.99 66.40 56.31 60.57 53.01 73.80 61.68
+Technological 61.49 64.96 56.12 62.00 50.60 74.21 61.56
+Avg. 60.03 63.24 55.48 62.03 50.96 74.09 60.97
 intfloat-e5-mistral-7b-instruct
-Biological 18.64 19.90 22.40 32.79 22.57 43.06 26.56
-Chemical 25.67 19.23 19.34 36.47 16.04 35.54 25.38
-Environmental 24.96 17.65 21.97 28.46 22.05 41.30 26.06
-Extraterrestrial 21.42 21.71 18.92 34.04 21.21 37.67 25.83
-Geohazard 21.45 17.30 16.98 25.16 19.57 41.11 23.59
-Meteorological&hydrological 20.67 15.84 20.54 24.92 20.85 37.62 23.41
-Societal 21.53 24.83 16.94 31.62 23.46 35.64 25.67
-Technological 18.86 19.62 18.73 34.85 17.34 44.61 25.67
-Avg. 21.65 19.51 19.48 31.04 20.39 39.57 25.27
+Biological 66.33 64.74 65.22 67.64 49.45 66.61 63.33
+Chemical 66.09 63.90 62.22 69.82 45.50 65.29 62.13
+Environmental 66.78 63.40 64.84 69.87 49.74 66.29 63.49
+Extraterrestrial 63.51 65.90 61.75 67.95 45.71 68.69 62.25
+Geohazard 65.17 63.03 61.52 66.19 45.90 63.45 60.88
+Meteorological&hydrological 66.28 65.67 63.71 67.20 47.76 65.58 62.70
+Societal 66.48 68.75 64.40 66.21 49.86 67.39 63.85
+Technological 64.58 65.91 63.73 68.63 47.54 67.85 63.04
+Avg. 65.65 65.16 63.42 67.94 47.68 66.39 62.71
 intfloat-e5-small-v2
-Biological 66.38 63.19 61.91 61.38 48.09 75.04 62.67
-Chemical 66.99 60.81 61.30 63.53 47.24 75.17 62.51
-Environmental 67.61 64.30 62.38 65.73 50.80 75.57 64.40
-Extraterrestrial 63.65 60.90 57.98 60.98 45.13 74.96 60.60
-Geohazard 65.73 61.10 59.06 60.21 43.82 74.60 60.75
-Meteorological&hydrological 66.89 63.88 61.80 62.53 46.50 74.76 62.73
-Societal 64.53 65.51 61.41 60.10 47.63 74.69 62.31
-Technological 65.19 64.31 61.32 61.88 47.51 73.84 62.34
-Avg. 65.87 63.00 60.89 62.04 47.09 74.83 62.29
+Biological 66.22 63.00 60.72 61.14 48.06 74.11 62.21
+Chemical 66.80 60.74 60.79 63.40 47.30 74.23 62.21
+Environmental 67.56 63.90 61.53 65.49 50.76 74.97 64.04
+Extraterrestrial 63.49 60.68 57.15 60.75 45.70 74.18 60.33
+Geohazard 65.60 60.94 58.51 59.69 44.13 73.69 60.42
+Meteorological&hydrological 66.63 63.81 61.09 62.35 46.44 73.46 62.30
+Societal 64.22 65.40 60.63 59.83 47.62 73.70 61.90
+Technological 64.74 64.24 60.35 61.59 46.94 73.10 61.83
+Avg. 65.66 62.84 60.10 61.78 47.12 73.93 61.90
 intfloat-multilingual-e5-large
-Biological 68.63 62.63 64.36 60.27 49.91 74.90 63.45
-Chemical 66.87 62.82 63.95 61.04 49.49 74.90 63.18
-Environmental 66.74 63.98 64.06 62.20 54.65 74.76 64.40
-Extraterrestrial 65.81 63.71 60.26 59.62 49.28 75.41 62.35
-Geohazard 66.65 62.97 63.71 58.47 49.68 75.10 62.76
-Meteorological&hydrological 68.57 63.87 63.41 60.45 50.40 74.53 63.54
-Societal 67.61 67.61 64.44 58.97 54.07 74.86 64.59
-Technological 67.47 66.39 65.80 61.08 50.71 76.02 64.58
-Avg. 67.29 64.25 63.75 60.26 51.02 75.06 63.61
+Biological 68.44 62.41 63.18 60.18 49.79 73.97 63.00
+Chemical 66.63 62.89 63.42 60.94 49.54 74.04 62.91
+Environmental 66.66 63.64 63.14 62.10 54.46 74.08 64.01
+Extraterrestrial 65.53 63.50 59.49 59.41 50.22 74.51 62.11
+Geohazard 66.68 62.72 63.21 58.09 50.76 74.09 62.59
+Meteorological&hydrological 68.26 63.78 62.80 60.25 50.22 73.31 63.10
+Societal 67.34 67.44 63.71 58.75 54.31 74.00 64.26
+Technological 67.13 66.28 64.96 60.76 50.31 75.16 64.10
+Avg. 67.08 64.08 62.99 60.06 51.20 74.14 63.26
 intfloat-multilingual-e5-large-instruct
-Biological 70.03 63.24 64.67 66.51 50.24 64.07 63.13
-Chemical 67.34 63.58 61.41 68.56 48.65 63.33 62.15
-Environmental 68.93 64.34 63.38 69.14 50.64 62.96 63.23
-Extraterrestrial 66.19 65.08 61.14 68.44 49.74 66.73 62.89
-Geohazard 68.16 62.03 62.75 64.45 46.34 62.08 60.97
-Meteorological&hydrological 68.59 66.05 63.42 66.52 50.51 64.22 63.22
-Societal 69.62 68.27 63.92 66.58 51.19 65.31 64.15
-Technological 68.25 66.61 65.26 67.51 47.72 65.81 63.53
-Avg. 68.39 64.90 63.24 67.21 49.38 64.31 62.91
+Biological 69.67 62.94 63.35 66.46 50.04 63.39 62.64
+Chemical 67.29 63.69 60.79 68.42 48.02 62.54 61.79
+Environmental 68.82 63.96 62.62 68.88 49.87 62.37 62.76
+Extraterrestrial 66.07 64.72 60.45 68.28 49.35 66.02 62.48
+Geohazard 67.91 61.88 62.11 63.99 45.83 61.20 60.49
+Meteorological&hydrological 68.24 65.91 62.79 66.38 49.79 63.12 62.70
+Societal 69.27 68.18 63.13 66.23 50.28 64.50 63.60
+Technological 67.85 66.46 64.43 67.08 46.77 65.11 62.95
+Avg. 68.14 64.72 62.46 66.96 48.75 63.53 62.43
 Table 25: Performance of evaluated models under six search intents and eight event types under the exact search setting. Part IV
 QA QAdoc Twitter FC NLI STS Avg.
-llmrails-ember-v1
-Biological 64.90 62.60 49.37 60.23 45.66 74.47 59.54
-Chemical 65.00 60.86 47.16 62.76 43.98 75.04 59.14
-Environmental 65.68 61.87 47.82 62.25 48.17 74.18 59.99
-Extraterrestrial 65.52 65.06 44.55 62.47 42.29 74.28 59.03
-Geohazard 63.42 60.96 44.08 59.80 43.70 72.78 57.46
-Meteorological&hydrological 64.02 64.12 45.55 60.74 42.42 72.11 58.16
-Societal 65.11 66.53 43.32 59.41 46.03 74.16 59.09
-Technological 64.87 64.73 45.61 60.78 44.23 74.90 59.19
-Avg. 64.82 63.34 45.93 61.06 44.56 73.99 58.95
+snowflake-arctic-embed-m-long
+Biological 19.40 10.57 21.85 39.39 41.02 52.71 30.82
+Chemical 19.97 10.58 21.06 36.59 40.40 55.39 30.66
+Environmental 21.57 11.81 21.01 34.99 48.55 55.06 32.16
+Extraterrestrial 21.22 11.89 17.32 35.89 41.66 56.66 30.77
+Geohazard 21.78 8.92 17.99 35.32 40.09 52.50 29.44
+Meteorological&hydrological 19.71 10.04 18.89 36.12 40.67 54.56 30.00
+Societal 23.30 11.95 20.69 35.64 43.05 56.58 31.87
+Technological 24.50 10.93 17.08 35.66 39.77 56.53 30.75
+Avg. 21.43 10.84 19.49 36.20 41.90 55.00 30.81
 mixedbread-ai-mxbai-embed-large-v1
-Biological 64.38 62.94 43.89 57.80 40.49 69.47 56.50
-Chemical 64.29 60.87 42.16 59.77 39.28 70.09 56.08
-Environmental 65.73 60.71 42.67 59.05 43.81 69.91 56.98
-Extraterrestrial 65.50 65.11 38.38 60.94 38.15 69.16 56.21
-Geohazard 62.01 60.48 40.11 56.35 40.05 67.06 54.34
-Meteorological&hydrological 64.65 62.91 39.82 57.32 38.05 67.04 54.97
-Societal 65.72 65.82 39.30 58.65 41.65 68.60 56.62
-Technological 64.26 64.83 38.31 57.85 40.30 69.16 55.78
-Avg. 64.57 62.96 40.58 58.47 40.22 68.81 55.93
+Biological 64.21 62.79 43.11 57.82 40.36 68.66 56.16
+Chemical 64.22 60.92 41.70 59.77 39.37 69.16 55.86
+Environmental 65.66 60.40 41.92 58.91 43.62 69.31 56.64
+Extraterrestrial 65.29 64.74 37.89 60.75 38.84 68.20 55.95
+Geohazard 61.90 60.32 39.84 55.92 40.44 66.14 54.09
+Meteorological&hydrological 64.24 62.74 39.44 57.19 37.94 66.12 54.61
+Societal 65.49 65.63 38.81 58.36 41.72 67.70 56.29
+Technological 63.97 64.75 37.88 57.70 39.80 68.42 55.42
+Avg. 64.37 62.79 40.07 58.30 40.26 67.96 55.63
 nvidia-NV-Embed-v1
-Biological 69.88 63.74 57.69 62.45 47.40 68.75 61.65
-Chemical 69.31 56.86 57.02 59.12 45.99 66.66 59.16
-Environmental 69.14 63.86 57.85 61.98 52.46 68.36 62.28
-Extraterrestrial 65.64 62.43 53.94 57.93 45.98 69.16 59.18
-Geohazard 66.26 62.29 55.90 58.56 48.33 65.33 59.44
-Meteorological&hydrological 69.42 63.50 58.28 61.66 47.69 68.60 61.52
-Societal 68.68 66.72 57.45 58.88 50.42 67.42 61.59
-Technological 68.59 64.78 56.57 59.77 48.24 68.62 61.10
-Avg. 68.36 63.02 56.84 60.04 48.31 67.86 60.74
+Biological 69.72 63.54 56.65 62.33 47.93 68.08 61.37
+Chemical 69.10 56.97 56.36 59.00 45.84 65.88 58.86
+Environmental 69.01 63.61 57.10 61.78 52.05 67.91 61.91
+Extraterrestrial 65.45 62.21 53.16 57.79 45.98 68.42 58.84
+Geohazard 66.10 62.08 55.43 58.16 48.53 64.47 59.13
+Meteorological&hydrological 69.13 63.37 57.63 61.52 47.61 67.50 61.13
+Societal 68.40 66.54 56.95 58.62 50.21 66.67 61.23
+Technological 68.21 64.63 55.75 59.57 47.84 67.94 60.66
+Avg. 68.14 62.87 56.13 59.85 48.25 67.11 60.39
 nvidia-NV-Embed-v2
-Biological 76.42 69.25 42.42 69.75 58.26 77.86 65.66
-Chemical 74.82 69.52 42.07 68.68 57.96 76.22 64.88
-Environmental 75.31 68.66 43.82 69.63 61.09 77.21 65.95
-Extraterrestrial 73.06 70.08 44.88 68.30 56.31 78.18 65.14
-Geohazard 73.91 67.35 41.35 66.72 57.88 75.91 63.85
-Meteorological&hydrological 75.33 69.91 43.11 68.37 58.88 76.85 65.41
-Societal 75.38 72.11 44.57 68.78 61.06 76.54 66.41
-Technological 73.91 71.01 43.18 68.85 58.42 77.31 65.45
-Avg. 74.77 69.74 43.18 68.64 58.73 77.01 65.34
-sentence-transformers-all-mpnet-base-v2
-Biological 14.00 7.64 17.18 43.65 27.58 35.38 24.24
-Chemical 13.09 9.93 14.53 48.81 24.14 35.84 24.39
-Environmental 12.45 10.01 18.17 47.32 30.32 35.48 25.62
-Extraterrestrial 16.24 8.11 14.82 48.79 27.48 42.02 26.24
-Geohazard 14.96 7.75 13.30 43.82 27.70 36.33 23.98
-Meteorological&hydrological 16.52 7.84 16.39 44.34 28.07 36.99 25.02
-Societal 20.10 14.49 18.41 45.98 27.31 39.06 27.56
-Technological 13.16 12.41 16.55 47.68 26.44 36.75 25.50
-Avg. 15.06 9.77 16.17 46.30 27.38 37.23 25.32
+Biological 76.22 69.00 41.41 69.62 58.10 77.03 65.23
+Chemical 74.61 69.41 41.58 68.47 57.64 75.33 64.51
+Environmental 75.20 68.39 43.19 69.37 60.63 76.54 65.55
+Extraterrestrial 72.83 69.73 44.26 68.16 56.37 77.37 64.79
+Geohazard 73.78 67.16 40.89 66.24 58.05 74.93 63.51
+Meteorological&hydrological 75.03 69.71 42.59 68.16 58.49 75.67 64.94
+Societal 75.06 71.88 44.08 68.51 60.55 75.63 65.95
+Technological 73.68 70.83 42.43 68.62 57.26 76.50 64.89
+Avg. 74.55 69.51 42.55 68.39 58.39 76.13 64.92
+thenlper-gte-small
+Biological 18.78 12.74 13.26 51.17 40.04 55.52 31.92
+Chemical 18.64 12.89 12.90 53.93 35.96 54.87 31.53
+Environmental 13.65 7.32 8.02 47.89 42.45 53.68 28.83
+Extraterrestrial 19.50 6.40 10.64 50.69 36.66 57.62 30.25
+Geohazard 16.16 5.30 9.30 45.85 36.16 53.58 27.72
+Meteorological&hydrological 17.88 8.92 11.34 49.99 37.82 55.61 30.26
+Societal 20.18 9.82 11.12 46.59 36.93 55.15 29.96
+Technological 20.84 9.26 11.74 52.41 35.69 57.71 31.27
+Avg. 18.20 9.08 11.04 49.81 37.71 55.47 30.22
 thenlper-gte-base
-Biological 10.09 5.90 45.04 61.49 44.70 49.42 36.11
-Chemical 9.73 5.62 40.67 62.27 41.06 53.75 35.52
-Environmental 6.03 4.75 38.62 63.00 48.95 40.21 33.59
-Extraterrestrial 10.42 4.54 33.58 59.45 41.39 44.64 32.34
-Geohazard 7.96 3.33 34.20 58.54 41.68 43.00 31.45
-Meteorological&hydrological 9.16 5.31 38.46 60.76 40.55 43.60 32.97
-Societal 10.00 7.88 39.76 60.39 44.40 48.17 35.10
-Technological 10.39 5.44 38.00 60.50 40.08 50.33 34.13
-Avg. 9.22 5.35 38.54 60.80 42.85 46.64 33.90
+Biological 9.95 5.91 44.29 61.38 44.74 48.79 35.84
+Chemical 9.69 5.57 40.17 62.04 41.18 53.04 35.28
+Environmental 6.01 4.73 38.02 62.79 48.67 39.85 33.35
+Extraterrestrial 10.36 4.50 33.16 59.23 41.91 44.01 32.19
+Geohazard 7.93 3.33 33.93 58.13 41.61 42.19 31.19
+Meteorological&hydrological 9.15 5.33 38.10 60.63 40.33 42.71 32.71
+Societal 9.95 7.88 39.41 60.21 44.04 47.51 34.83
+Technological 10.22 5.44 37.40 60.22 39.95 49.82 33.84
+Avg. 9.16 5.34 38.06 60.58 42.80 45.99 33.65
 Table 26: Performance of evaluated models under six search intents and eight event types under the exact search setting. Part V